### v1

# MCbiF: Measuring Topological Autocorrelation in Multiscale Clusterings via 2-Parameter Persistent Homology

Abstract
Datasets often possess an intrinsic multiscale structure with meaningful descriptions at different levels of coarseness. Such datasets are naturally described as multi-resolution clusterings, i.e., not necessarily hierarchical sequences of partitions across scales. To analyse and compare such sequences, we use tools from topological data analysis and define the Multiscale Clustering Bifiltration (MCbiF), a 2-parameter filtration of abstract simplicial complexes that encodes cluster intersection patterns across scales. The MCbiF can be interpreted as a higher-order extension of Sankey diagrams and reduces to a dendrogram for hierarchical sequences. We show that the multiparameter persistent homology (MPH) of the MCbiF yields a finitely presented and block decomposable module, and its stable Hilbert functions characterise the topological autocorrelation of the sequence of partitions. In particular, at dimension zero, the MPH captures violations of the refinement order of partitions, whereas at dimension one, the MPH captures higher-order inconsistencies between clusters across scales. We demonstrate through experiments the use of MCbiF Hilbert functions as topological feature maps for downstream machine learning tasks. MCbiF feature maps outperform information-based baseline features on both regression and classification tasks on synthetic sets of non-hierarchical sequences of partitions. We also show an application of MCbiF to real-world data to measure non-hierarchies in wild mice social grouping patterns across time.
Keywords: topological data analysis, multiparameter persistent homology, multiscale clustering, non-hierarchical clustering, Sankey diagrams
1 Introduction
In many applications, datasets possess an intrinsic multiscale structure, whereby meaningful descriptions exist at different scales, i.e., at different resolutions or levels of coarseness. Think, for instance, of the multi-resolution structure in commuter mobility patterns [Alessandretti et al., 2020, Schindler et al., 2023], communities in social networks [Beguerisse-Díaz et al., 2017] and thematic groups of documents [Blei et al., 2003, Grootendorst, 2022]; the subgroupings in single-cell data [Hoekzema et al., 2022] or phylogenetic trees [Chan et al., 2013]; and the functional substructures in proteins [Delvenne et al., 2010, Delmotte et al., 2011]. In such cases, the natural description of the dataset goes beyond a single clustering and consists of a multi-resolution sequence of partitions across scales parametrised by a scale parameter $t$. Traditionally, multiscale descriptions have emerged from hierarchical clustering, where $t$ corresponds to the depth of the dendrogram [Carlsson and Mémoli, 2010, Murtagh and Contreras, 2012]. However, in many important real-world applications, the data structure is multiscale, yet non-hierarchical. Examples include temporal clustering, where $t$ corresponds to physical time [Rosvall and Bergstrom, 2010, Liechti and Bonhoeffer, 2020, Bovet et al., 2022]; topic modelling and document classification, where $t$ captures the coarseness of the topic groupings [Altuncu et al., 2019, Fukuyama et al., 2023, Liu et al., 2025]; and generic multiscale clusterings for data that result from exploiting a diffusion on the data geometry, where $t$ is the increasing time horizon of the diffusion [Coifman et al., 2005, Azran and Ghahramani, 2006, Lambiotte et al., 2014].
A natural problem is then how to analyse and compare non-hierarchical multi-resolution sequences of partitions that are organised by the scale $t$. Here we address this question from the perspective of topological data analysis [Carlsson and Zomorodian, 2009, Carlsson et al., 2009, Botnan and Lesnick, 2023] by introducing the Multiscale Clustering Bifiltration (MCbiF), a 2-parameter filtration of abstract simplicial complexes that encodes the patterns of cluster intersections across all scales.
Problem definition.
A partition $\pi$ of a finite set $X=\{x_{1},x_{2},...,x_{N}\}$ is a collection of mutually exclusive subsets $C_{i}\subseteq X$ (here called clusters) that cover $X$, i.e., $\pi=\{C_{1},\dots,C_{c}\}$ such that $X=~\bigcup_{i=1}^{c}C_{i}$, and $C_{i}\bigcap C_{j}=\emptyset,\,\forall i\neq j$. The cardinality $|\pi|=c$ is the number of clusters in $\pi$ and, for notational convenience, we use $\pi_{i}$ to denote the $i$-th cluster $C_{i}$ of $\pi$.
Let $\Pi_{X}$ denote the space of partitions of $X$. We write $\pi\leq\pi^{\prime}$ if every cluster in $\pi$ is contained in a cluster of $\pi^{\prime}$. This refinement relation constitutes a partial order and leads to the partition lattice $(\Pi_{X},\leq)$ with lower bound $\hat{0}:=\{\{x_{1}\},\dots,\{x_{N}\}\}$ and upper bound $\hat{1}:=\{X\}$ [Birkhoff, 1967].
Here, we consider a sequence of partitions defined as a piecewise-constant function $\theta:[t_{1},\infty)\rightarrow\Pi_{X},\,t\mapsto\theta(t)\in\Pi_{X}$ such that a partition of $X$ is assigned to each $t$, and the scale index $t\in[t_{1},\infty)$ has $M$ change points $t_{1}<t_{2}<...<t_{M}$. In particular, $\theta(t)=\theta(t_{m})$ for $t\in[t_{m},t_{m+1})$, $m=1,\dots,M-1$, and $\theta(t)=\theta(t_{M})$ for $t\in[t_{M},\infty)$. The sequence $\theta$ is called hierarchical if either $\theta(s)\leq\theta(t),\,\forall s\leq t$, or $\theta(s)\geq\theta(t),\,\forall s\leq t$. The sequence $\theta$ is called coarse-graining if $|\theta(s)|\geq|\theta(t)|,\,\forall s\leq t$.  [^1] Coarse-graining is equivalent to non-decreasing mean cluster size (see Remark 37 in Appendix A.1). Conversely, $\theta$ is called fine-graining if $|\theta(s)|\leq|\theta(t)|,\,\forall s\leq t$.
Our goal is to characterise and analyse arbitrary sequences of partitions $\theta$, including non-hierarchical ones, in an integrated manner, taking account of memory effects across the scale $t$.
Remark 1.
Here, we are not concerned with the task of computing the multiscale clustering (i.e., the sequence of partitions $\theta$) from dataset $X$, for which several methods exist. Rather, we take $\theta$ as a given, and we aim to analyse its structure.
Remark 2.
This problem is distinct from consensus clustering, which aims to produce a summary partition by combining a set of partitions obtained, e.g., from different optimisations or clustering algorithms [Strehl and Ghosh, 2002, Vega-Pons and Ruiz-Shulcloper, 2011].
Figure 1: (a) Illustration of how the MCbiF encodes the structure of a non-hierarchical sequence of partitions $\theta$ as a bifiltration of abstract simplicial complexes $K^{s,t}$. See Example 42 for a detailed description. (b) The Hilbert functions $\mathrm{HF}_{k}(s,t)$ of the MCbiF are invariants that capture the topological autocorrelation of $\theta$: violations of the refinement order at dimension $k=0$, and higher-order cluster inconsistencies at dimension $k=1$. The Hilbert functions can be used as feature maps for downstream machine learning tasks.
Contributions.
To address this problem, we define the MCbiF, a bifiltration of abstract simplicial complexes, which represents the clusters and their intersection patterns in the sequence $\theta$ for varying starting scale $s$ and lag tst-s (Fig. 1). Using the machinery of multiparameter persistent homology (MPH) [Carlsson and Zomorodian, 2009, Carlsson et al., 2009, Botnan and Lesnick, 2023], we prove that the MCbiF leads to a block decomposable persistence module with stable Hilbert functions $\mathrm{HF}_{k}(s,t)$, and we show that these invariants serve as measures of the topological autocorrelation of the sequence of partitions $\theta$ across the scale $t$. In particular, $\mathrm{HF}_{k}(s,t)$ quantifies the non-hierarchy in $\theta$ in two complementary ways: at dimension $k=0$, it detects the lack of a maximal partition in the subposet $\theta([s,t])$ with respect to refinement, and at dimension $k=1$, it quantifies the higher-order inconsistencies of cluster assignments across scales. In contrast, baseline methods such as ultrametrics [Carlsson and Mémoli, 2010] or information-based measures [Meilă, 2003] are restricted to pairwise comparisons between, respectively, elements or clusters; hence these methods cannot detect higher-order cluster inconsistencies. Furthermore, we provide an equivalent nerve-based construction of the MCbiF that can be interpreted as a higher-order extension of the Sankey diagram of the sequence of partitions. In the hierarchical case, the 1-dimensional MPH of the MCbiF is trivial, and the Sankey diagram reduces to a dendrogram such that the 0-dimensional MCbiF Hilbert function can be obtained from the number of branches in the dendrogram.
The Hilbert functions of the MCbiF provide interpretable feature maps that can be used in downstream machine learning tasks. In our experiments, the MCbiF feature maps outperform information-based baseline features [Meilă, 2007] on both regression and classification tasks on non-hierarchical sequences of partitions. We also show an application of MCbiF to real-world data to measure non-hierarchies in wild mice social grouping patterns across time [Bovet et al., 2022].
2 Related Work
Dendrograms and Ultrametrics.
A hierarchical, coarse-graining sequence $\theta$ with $\theta(t_{1}=0)=\hat{0}$ and $\theta(t_{M})=\hat{1}$ is called an agglomerative dendrogram, and can be represented by an acyclic rooted merge tree [Jain et al., 1999, Carlsson and Mémoli, 2010]. One can define an ultrametric $D_{\theta}$ from the first-merge times, which corresponds to the depth in the dendrogram. Carlsson and Mémoli [2010] showed that there is a one-to-one correspondence between agglomerative dendrograms and ultrametrics, which can be used to efficiently compare two such dendrograms via the Gromov-Hausdorff distance between the ultrametric spaces [Mémoli et al., 2023]. When $\theta$ is non-hierarchical, however, the first-merge times no longer define the sequence uniquely because clusters that have merged can split off again. In this case, $\theta$ cannot be represented by a tree and $D_{\theta}$ does not fulfil the triangle inequality in general. Hence, ultrametrics cannot be used to analyse and compare non-hierarchical sequences of partitions (see Section 5).
Information-based Comparison of Clusterings.
Information-based measures can be used to compare a pair of partitions. Assuming a uniform distribution on $X$, one can derive probability distributions for partitions interpreted as random variables and thus measure the information gain and loss between two partitions using the conditional entropy (CE) or the variation of information (VI), which is a metric on $\Pi_{X}$ [Meilă, 2003, 2007]. See Appendix E for detailed formulas. Extending information-based measures to more than two partitions is non-trivial. In consensus clustering, the average VI is used as a consensus index (CI) [Vinh and Epps, 2009, Vinh et al., 2010] for multiple partitions. However, the CI is independent of the ordering in the sequence and so cannot capture memory effects in sequences of partitions. Another limitation of these measures is that they rely only on the joint probability between pairs of random variables, hence higher-order cluster inconsistencies are not captured (see Section 5).
3 Background
3.1 Sankey Diagrams
Non-hierarchical sequences of partitions $\theta$ are visualised by $M$-layered flow graphs $S(\theta)=(V=V_{1}\uplus...\uplus V_{M},E=E_{1}\uplus...\uplus E_{M-1})$ called Sankey diagrams [Sankey, 1898, Zarate et al., 2018], where each level $m=1,\ldots,M$ corresponds to a partition and vertices $V_{m}$ represent its clusters while the directed edges $E_{m}$ between levels indicate the overlap between clusters:
$V_{m}:=\{(m,i)\;|\;1\leq i\leq|\theta(t_{m})|\}\quad\text{and}\quad E_{m}=\left\{\left[(m,i),(m+1,j)\right]\;|\;\theta(t_{m})_{i}\cap\theta(t_{m+1})_{j}\neq\emptyset\right\},$ (1)
where $u,v]\in E_{m}$ denotes a directed edge from $u\in V_{m}$ to $v\in V_{m+1}$. If $\theta$ is hierarchical, the Sankey diagram $S(\theta)$ is a directed tree—a merge-tree if $\theta$ is coarse-graining, or a split-tree if $\theta$ is fine-graining. The graph $S(\theta)$ is sometimes also called an alluvial diagram [Rosvall and Bergstrom, 2010].
Sankey diagrams are studied in computer graphics as they allow for the visualisation of complex relational data. In this context, a Sankey diagram is represented as a layout on the plane, whereby the nodes in each layer $V_{m}$ are vertically ordered according to a ranking $\tau_{m}:V_{m}\rightarrow\{1,\dots,|V_{m}|\}$, and the layout of the Sankey diagram is then defined by the collection of such rankings, $\tau:=(\tau_{1},\dots,\tau_{M})$. For visualisation purposes, the layered layout should ideally minimise the number of crossings between consecutive layers, where a crossing between two edges $u,v],[u^{\prime},v^{\prime}]\in E_{m}$ occurs if $\tau_{m}(u)>\tau_{m}(u^{\prime})$ and $\tau_{m+1}(v)<\tau_{m+1}(v^{\prime})$ or vice versa, and the crossing number [Warfield, 1977] is given by:
$\kappa_{\theta}(\tau):=\sum_{m=1}^{M-1}\sum_{[u,v],[u^{\prime},v^{\prime}]\in E_{m}}\mathbbm{1}_{\tau_{m}(u)>\tau_{m}(u^{\prime})\land\tau_{m+1}(v)<\tau_{m+1}(v^{\prime})},$ (2)
where $\mathbbm{1}$ denotes the indicator function. The crossing number $\kappa_{\theta}(\tau)$ of the layout of the Sankey diagram $S(\theta)$ can be minimised by permuting the rankings in the layers, $\tau_{m}$, and we denote the minimum crossing number for the layout as:
$\overline{\kappa}_{\theta}:=\min_{\tau}\kappa_{\theta}(\tau).$ (3)
This problem is known to be NP-complete [Garey and Johnson, 2006] and finding efficient optimisation algorithms is an active research area [Zarate et al., 2018, Li et al., 2025].
3.2 Multiparameter Persistent Homology
Multiparameter persistent homology (MPH) is an extension of standard persistent homology to $n>1$ parameters, first introduced by Carlsson and Zomorodian [2009]. We present here basic definitions, see Carlsson and Zomorodian [2009], Carlsson et al. [2009], Botnan and Lesnick [2023] for details.
Simplicial Complex.
Let $K$ be a simplicial complex defined for the set $X$, such that $K\subseteq 2^{X}$ and $K$ is closed under the operation of building subsets. The elements of $\sigma\in K$ are called simplices and a $k$-dimensional simplex (or $k$-simplex) can be represented as $\sigma=[x_{1},...,x_{k+1}]$ where $x_{1},\dots,x_{k+1}\in X$ and we have fixed an arbitrary order on $X$. Note that $k=0$ corresponds to vertices, $k=1$ to edges, and $k=2$ to triangles. We define the $k$-skeleton $K_{k}$ of $K$ as the union of its $n$-simplices for $n\leq k$. We also define $\dim(K)$ as the largest dimension of any simplex in $K$.
Multiparameter Filtration.
Let us define the parameter space $(P,\leq)$ as the product of $n\geq 1$ partially ordered sets $P=P_{1}\times\dots\times P_{n}$, i.e., $\boldsymbol{a}\leq\boldsymbol{b}$ for $\boldsymbol{a},\boldsymbol{b}\in P$ if and only if $\boldsymbol{a}_{i}\leq\boldsymbol{b}_{i}$ in $P_{i}$ for $i=1,\dots,n$. A collection of subcomplexes $(K^{\boldsymbol{a}})_{\boldsymbol{a}\in\mathbb{R}^{n}}$ with $K=\bigcup_{\boldsymbol{a}\in\mathbb{R}^{n}}K^{\boldsymbol{a}}$ and inclusion maps $\{i_{\boldsymbol{a},\boldsymbol{b}}:K^{\boldsymbol{a}}\rightarrow K^{\boldsymbol{b}}\}_{\boldsymbol{a}\leq\boldsymbol{b}}$ that yield a commutative diagram is called a multiparameter filtration (or bifiltration for $n=2$). We denote by $\mathrm{birth}(\sigma)\subseteq P$ the set of parameters, called multigrades (or bigrades for $n=2$), at which simplex $\sigma\in K$ first appears in the filtration. For example, the sublevel filtration $K^{\boldsymbol{a}}=\{\sigma\in K\;|\;f(\sigma)\leq\boldsymbol{a}\}$ for a filtration function $f:K\rightarrow P$ maps each simplex $\sigma$ to a unique multigrade $f(\sigma)$, i.e., $|\mathrm{birth}(\sigma)|=1$. A filtration is called one-critical if it is isomorphic to a sublevel filtration, and multi-critical otherwise.
Multiparameter Persistent Homology.
Let $H_{k}$ for $k\in\{0,\dots,\dim(K)\}$ denote the $k$-dimensional homology functor with coefficients in a field [Hatcher, 2002], see Appendix C for details. Then $H_{k}$ applied to the multiparameter filtration leads to a multiparameter persistence module, i.e., a collection of vector spaces $(H_{k}(K^{\boldsymbol{a}}))_{\boldsymbol{a}\in\mathbb{R}^{n}}$, which are the homology groups whose elements are the generators of $k$-dimensional non-bounding cycles, and linear maps $\{\imath_{\boldsymbol{a},\boldsymbol{b}}:=H_{k}(i_{\boldsymbol{a},\boldsymbol{b}}):H_{k}(K^{\boldsymbol{a}})\rightarrow H_{k}(K^{\boldsymbol{b}})\}_{\boldsymbol{a}\leq\boldsymbol{b}}$ that yield a commutative diagram called multiparameter persistent homology (MPH). For dimension $k=0$, $H_{k}$ captures the number of disconnected components and for $k=1$, the number of holes. Note that, for $n=1$, we recover standard persistent homology (PH)  [Edelsbrunner et al., 2002].
Hilbert Function.
While barcodes are complete invariants of 1-parameter PH ($n=1$), the more complicated algebraic structure of MPH ($n\geq 2$) does not allow for such simple invariants in general; hence, various non-complete invariants of the MPH are used in practice. We focus on the $k$-dimensional Hilbert function [Botnan and Lesnick, 2023] defined as
$\mathrm{HF}_{k}:P\rightarrow\mathbb{N}_{0},\,\mathbf{a}\mapsto\operatorname{rank}[H_{k}(i_{\boldsymbol{a},\boldsymbol{a}})]=\dim[H_{k}(K^{\mathbf{a}})],$ (4)
which maps each filtration index $\mathbf{a}$ to the $k$-dimensional Betti number of the corresponding complex $K^{\mathbf{a}}$. We call the $k$-dimensional MPH trivial if $\mathrm{HF}_{k}=0$. The Hilbert distance is then defined as the $L_{2}$ norm on the space of Hilbert functions and can be used to compare multiparameter persistence modules.
4 The Multiscale Clustering Bifiltration (MCbiF)
The central object of our paper is a novel bifiltration of abstract simplicial complexes that encodes cluster intersection patterns in the sequence of partitions $\theta$ across the scale $t$.
Definition 3 (Multiscale Clustering Bifiltration).
Let $\theta:[t_{1},\infty)\rightarrow\Pi_{X}$ be a sequence of partitions. We define $\mathcal{M}$, the Multiscale Clustering Bifiltration (MCbiF) as a bifiltration of abstract simplicial complexes:
$\mathcal{M}:=(K^{s,t})_{t_{1}\leq s\leq t}\quad\text{where}\quad K^{s,t}:=\bigcup_{t_{1}\leq s\leq r\leq t}\,\,\bigcup_{C\in\theta(r)}\Delta C,\quad t_{1}\leq s\leq t.$ (5)
In this construction, each cluster $C$ corresponds to a (|C|1)(|C|-1)-dimensional solid simplex $\Delta C:=2^{C}$, which, by definition, contains all its lower dimensional simplices [Schindler and Barahona, 2025]. This echoes natural concepts of data clustering as information compression or lumping [Rosvall and Bergstrom, 2008, 2011, Lambiotte et al., 2014], and of clusters as equivalence classes [Brualdi, 2010]. The MCbiF then aggregates all clusters (simplices) from partition $\theta(s)$ to $\theta(t)$ through the union operators, such that a $k$-simplex $\sigma=[x_{1},\dots,x_{k+1}]\in K^{s,t}$ consists of elements that are assigned to the same cluster (at least once) in the interval $[s,t]$, i.e., $x_{1},\dots,x_{k+1}\in C$ for some cluster $C\in\theta(r)$ and $r\in[s,t]$. The bifiltration thus depends not only on the lag |ts||t-s| but also on the starting scale $s$, and captures the topological autocorrelation in the sequence of partitions, see Fig. 1.
We first show that the MCbiF is indeed a well-defined bifiltration.
Proposition 4.
$\mathcal{M}$ is a multi-critical bifiltration uniquely defined by its values on the finite grid $P=\{(s,t)\in[t_{1},\dots,t_{M}]\times[t_{1},\dots,t_{M}]\;|\;s\leq t\}$ with partial order $(s,t)\leq(s^{\prime},t^{\prime})$ if $s\geq s^{\prime},t\leq t^{\prime}$.
The proof is straightforward and can be found in Appendix A.2.
The MCbiF leads to a triangular commutative diagram where the arrows indicate inclusion maps between abstract simplicial complexes (see Fig. 1). The sequence of partitions $\theta(t)$ is encoded by the complexes $K^{t,t}$ on the diagonal of the diagram. Moving along horizontal arrows corresponds to fixing a starting scale $s$ and going forward in the sequence $\theta$, thus capturing the coarse-graining of partitions. Moving along vertical arrows corresponds to fixing an end scale $t$ and aggregating $\theta$ going backwards, capturing the fine-graining of partitions.
Remark 5.
By fixing $s:=t_{1}$ (i.e., the top row in the commutative MCbiF diagram), we recover the 1-parameter Multiscale Clustering Filtration (MCF) defined by Schindler and Barahona [2025]. MCF was designed to quantify non-hierarchies in coarse-graining sequences of partitions and thus cannot capture fine-graining. For example, a large cluster $C\in\theta(s^{\prime})$ prevents MCF from detecting cluster assignment conflicts between elements $x,y\in C$ for $t\geq s^{\prime}$, see Section 4.1. In contrast, MCbiF encodes the full topological autocorrelation contained in $\theta$ by varying both the starting scale $s$ and the lag tst-s.
Applying MPH to the bifiltration $\mathcal{M}$ at dimensions $k\leq\dim K$, for $K=K^{t_{M},t_{M}}$, leads to a triangular diagram of simplicial complexes $H_{k}(K^{s,t})$ called persistence module (see Section 3.2). This persistence module has strong algebraic properties, as stated in the following proposition.
Proposition 6.
For any $k\leq\dim K$, the MCbiF persistence module $H_{k}(K^{s,t})$ is pointwise finite-dimensional, finitely presented and block-decomposable.
See Botnan and Lesnick [2023] for definitions and Appendix A.2 for details and a full proof.
The properties in Proposition 6 are important because they guarantee algebraic stability of the MCbiF [Bjerkevik, 2021]. In particular, the finite presentation property implies stability of the MCbiF Hilbert functions $\mathrm{HF}_{k}(s,t)$ (Eq. 4) with respect to small changes in the module [Oudot and Scoccola, 2024, Corollary 8.2.]. This justifies the use of $\mathrm{HF}_{k}(s,t)$ as simple interpretable invariants for the topological autocorrelation captured by MCbiF, as exploited in Section 4.1.
4.1 Measuring Topological Autocorrelation with MCbiF
We now show how topological autocorrelation as measured by $\mathrm{HF}_{k}(s,t)$ can be used to detect cluster-assignment conflicts. We focus on dimensions $k=0,1$, for which MPH is implemented in RIVET [Lesnick and Wright, 2015]. Note that $\mathrm{HF}_{0}(s,t)$ counts the number of connected components and $\mathrm{HF}_{1}(s,t)$ the number of 1-dimensional holes in $K^{s,t}$, see Section 3.2 for details. We show below that the computation of these invariants reveals different aspects of the non-hierarchy in the sequence of partitions.
4.1.1 Low-order Non-Hierarchy in Sequences of Partitions
Hierarchy in a sequence of partitions can be understood as a refinement of partitions captured by the partition lattice.
Definition 7 (Hierarchy).
The sequence of partitions $\theta$ is hierarchical in $[s,t]$ if we have a strict sequence of refinements: either $\theta(r_{1})\leq\theta(r_{2}),\forall r_{1},r_{2}\in[s,t]$ with $r_{1}\leq r_{2}$ (agglomerative) or $\theta(r_{1})\geq\theta(r_{2}),\,\forall r_{1},r_{2}\in[s,t]$ with $r_{1}\leq r_{2}$ (divisive). We say that $\theta$ is strictly hierarchical if it is hierarchical in $t_{1},\infty)$.
One important aspect of hierarchy is the nestedness of the clusters in the sequence.
Definition 8 (Nestedness).
We say that $\theta$ is nested in $[s,t]$ when $\forall r_{1},r_{2}\in[s,t]$, we have that $\forall C\in\theta(r_{1}),C^{\prime}\in\theta(r_{2})$, one of the sets $\setminus C^{\prime}, CC$$^{\prime}\setminus C$ or $C\cap C^{\prime}$ is empty. See [Korte and Vygen, 2012, Definition 2.12]. We say that $\theta$ is strictly nested when $\theta$ is nested in $t_{1},\infty)$.
Remark 9.
It follows directly from the definitions that a hierarchical sequence $\theta$ is always nested. However, nestedness does not necessarily imply hierarchy, as illustrated by the example in Fig 3b.
We can quantify the low-order non-hierarchy in a sequence $\theta$ through the computation of the invariant $\mathrm{HF}_{0}(s,t)$ and the associated notion of 0-conflicts defined next.
Remark 10.
Each partition $\theta(t)$ can be interpreted as an equivalence relation $\sim_{t}$ given by the property of belonging to the same cluster, i.e., $x\sim_{t}y$ if $\exists\;C\in\theta(t)$ such that $x,y\in C$ [Brualdi, 2010].
Definition 11 (0-conflict and triangle 0-conflict).
  1. a)
    We say that $\theta$ has a 0-conflict in $[s,t]$ if the subposet $\theta([s,t])$ has no maximum, i.e., $\nexists r\in[s,t]$ such that $\theta(r^{\prime})\leq\theta(r)$, $\forall r^{\prime}\in[s,t]$.
  2. b)
    We say that $\theta$ has a triangle 0-conflict in $[s,t]$ if $\exists\;x,y,z\in X$ such that $\exists r_{1},r_{2}\in[s,t]$: $x\sim_{r_{1}}y\sim_{r_{2}}z$ and $\nexists r\in[s,t]$: $x\nsim_{r}y\nsim_{r}z$.
Next, we show that all triangle 0-conflicts are also 0-conflicts. Moreover, all 0-conflicts break hierarchy and triangle 0-conflicts additionally break nestedness.
Proposition 12.
  1. (i)
    Every triangle 0-conflict is a 0-conflict, but the opposite is not true.
  2. (ii)
    If $\theta$ has a 0-conflict in $[s,t]$, then $\theta$ is non-hierarchical in $[s,t]$.
  3. (iii)
    If $\theta$ is either coarse- or fine-graining but non-hierarchical in $[s,t]$, then $\theta$ has a 0-conflict in $[s,t]$.
  4. (iv)
    If $\theta$ has a triangle 0-conflict in $[s,t]$, then $\theta$ is non-nested in $[s,t]$.
See Appendix A.2.1 for the simple proof. Fig. 3b illustrates a 0-conflict that is not a triangle 0-conflict, and Fig. 3c shows a triangle 0-conflict.
Remark 13.
Non-nestedness and non-hierarchy do not imply the presence of a 0-conflict. To see this, consider the simple counter-example given by $\theta(0)=\{\{x,y\},\{z\}\}$, $\theta(1)=\hat{1}$, $\theta(2)=\{\{x\},\{y,z\}\}$, which is non-nested but the partition $\theta(1)$ is the maximum of the subposet $\theta([0,1,2])$. This illustrates the need for the additional assumption of coarse- or fine-graining of $\theta$ in Proposition 12 (iii) for the condition of no 0-conflict to imply hierarchy.
The following proposition develops a sharp upper bound for $\mathrm{HF}_{0}$ that can be used to capture 0-conflicts.
Proposition 14.
  1. (i)
    $\mathrm{HF}_{0}(s,t)\leq\min_{r\in[s,t]}|\theta(r)|$, $\forall[s,t]\subseteq[t_{1},\infty)$.
  2. (ii)
    $\mathrm{HF}_{0}(s,t)<\min_{r\in[s,t]}|\theta(r)|$ iff $\theta$ has a 0-conflict in $[s,t]$.
  3. (iii)
    $\mathrm{HF}_{0}(s,t)=|\theta(r)|$ for $r\in[s,t]$ iff $\theta(r)$ is the maximum of the subposet $\theta([s,t])$.
See Appendix A.2.1 for a full proof. Proposition 14 shows that $\mathrm{HF}_{0}$ measures low-order non-hierarchy in $\theta$ by capturing 0-conflicts. To quantify this, we introduce a global normalised measure for the sequence $\theta$, defined as follows.
Definition 15 (Average 0-conflict).
Let $T:=t_{M}+\frac{t_{M}-t_{1}}{M-1}$. The average 0-conflict is defined as:
$0\,\leq\,\bar{c}_{0}(\theta):=1-\frac{2}{|T-t_{1}|^{2}}\int_{t_{1}}^{T}\int_{s}^{T}\frac{\mathrm{HF}_{0}(s,t)}{\min_{r\in[s,t]}\mathrm{HF}_{0}(r,r)}\mathrm{d}s\;\mathrm{d}t\,\leq\,1.$ (6)
Higher values of $\bar{c}_{0}(\theta)$ indicate a high level of 0-conflicts and increased low-order non-hierarchy, as shown by the next corollary.
Corollary 16.
  1. (i)
    If $\theta$ is hierarchical in $[s,t]$, then $\mathrm{HF}_{0}(s,t)=\min(|\theta(s)|,|\theta(t)|)$. As a special case, this implies $\mathrm{HF}_{0}(t,t)=|\theta(t)|,\forall t\geq t_{1}$.
  2. (ii)
    $\bar{c}_{0}(\theta)>0$ iff $\theta$ has a 0-conflict.
  3. (iii)
    Let $\theta$ be either coarse- or fine-graining. Then, $\bar{c}_{0}(\theta)=0$ iff $\theta$ is strictly hierarchical.
A proof can be found in Appendix A.2.1.
We can further detect triangle 0-conflicts by analysing the graph-theoretic properties of the MCbiF 1-skeleton $K_{1}^{s,t}$. Recall that the clustering coefficient $\mathrm{C}$ of a graph is defined as the ratio of the number of triangles to the number of paths of length 2 in the graph  [Luce and Perry, 1949, Newman, 2018].
Proposition 17.
$\mathrm{C}(K^{s,t}_{1})<1$ iff there is a triple $x,y,z\in X$ that leads to a triangle 0-conflict for $[s,t]$, and which is not a cycle, i.e., additionally to property b) in Definition 11 we also have $\nexists r_{3}\in[s,t]$: $x\sim_{r_{3}}z$.
See Appendix A.2.1 for a proof. Let us consider the graph generated as the disjoint union of all clusters from partitions in [s,t] as cliques. This graph is equivalent to the MCbiF 1-skeleton $K_{1}^{s,t}$. Proposition 17 shows that the clustering coefficient of this graph can be used to detect triangle 0-conflicts that are not cycles. To be able to detect triangle 0-conflicts that correspond to non-bounding cycles, we turn to the 1-dimensional homology in the next section.
4.1.2 Higher-order Inconsistencies between Clusters in Sequences of Partitions
Measuring 0-conflicts in $\theta$ is only one way of capturing non-hierarchy. An additional phenomenon that can arise in non-hierarchical sequences is higher-order inconsistencies of cluster assignments across scales. These are captured by the 1-dimensional homology groups [Schindler and Barahona, 2025] and the associated notion of 1-conflict, which we define next.
Recall the definition of $1$-cycles $Z_{1}(K^{s,t})$ and non-bounding cycles $H_{1}(K^{s,t})$ (Eq.14) summarised in Appendix C.
Definition 18 (1-conflict).
We say that $\theta$ has a 1-conflict in $[s,t]$ if $\exists\;x_{1},\dots,x_{n}\in X$ such that the 1-cycle $z=[x_{1},x_{2}]+\dots+[x_{n-1},x_{n}]+[x_{n},x_{1}]\in Z_{1}(K^{s,t})$ is non-bounding; in other words, $z]\in H_{1}(K^{s,t})$ with $z]\neq 0$.
The number of distinct 1-conflicts for the interval $[s,t]$ (up to equivalence of the homology classes) is given by $\mathrm{HF}_{1}(s,t)$. We first show that 1-conflicts also lead to triangle 0-conflicts and thus break hierarchy and nestedness of $\theta$.
Proposition 19.
  1. (i)
    $\mathrm{HF}_{1}(s,t)\geq 1$ iff $\theta$ has a 1-conflict in $[s,t]$.
  2. (ii)
    If $\theta$ has a 1-conflict in $[s,t]$, then it also has a triangle 0-conflict.
  3. (iii)
    If $\theta$ is hierarchical in $[s,t]$, then $\mathrm{HF}_{1}(s,t)=0$.
See Appendix A.2.1 for a proof. Proposition 19 shows that a 1-conflict is a special kind of triangle 0-conflict arising from higher-order cluster inconsistencies across scales. This is illustrated in Fig. 12 and more details are provided in Examples 42 and 43 in Appendix B, which present sequences of partitions where different 1-conflicts emerge across scales. Moreover, Proposition 19 (iii) shows that the MCbiF has a trivial 1-dimensional MPH if $\theta$ is strictly hierarchical.
Remark 20.
Proposition 19 states that every 1-conflict is a triangle 0-conflict. However, not every (triangle) 0-conflict is a 1-conflict, see Example 42. Note also that several triangle 0-conflicts in the sequence $\theta$ can lead to a 1-conflict, when the triangle 0-conflicts are linked together in such a way as to form a non-bounding cycle, see Example 42. We can test for these systematically using $\mathrm{HF}_{1}$.
Remark 21.
The presence of a 1-conflict for the interval $[s,t]$ signals the fact that assigning all the elements involved in the conflict to a shared cluster would increase the consistency of the sequence. Hence, when a 1-conflict gets resolved, e.g., the corresponding homology generator dies in the MPH at $(s,t^{\prime})$, $t<t^{\prime}$, then we say that $\theta(t^{\prime})$ is a conflict-resolving partition, see Schindler and Barahona [2025]. As two illustrations, consider Example 42, where a 1-conflict emerges through the interplay of partitions at scales $t=1,2,3$ and gets resolved by $\theta(4)=\hat{1}$, and Example 43, where three different 1-conflicts first emerge between scales $t=1$ and $t=6$ and then get resolved one by one by partitions $\theta(7)$, $\theta(8)$ and $\theta(9)$.
To quantify the presence of 1-conflicts in $\theta$ we introduce an unnormalised global measure for the sequence $\theta$, as follows.
Definition 22 (Average 1-conflict).
Let $T$ be defined as in Definition 15. The average 1-conflict is defined as:
$0\leq\bar{c}_{1}(\theta):=\frac{2}{|T-t_{1}|^{2}}\int_{t_{1}}^{T}\int_{s}^{T}\mathrm{HF}_{1}(s,t)\mathrm{d}s\;\mathrm{d}t.$ (7)
Corollary 23.
$c_{1}(\theta)>0$ iff $\theta$ has a 1-conflict. In particular, if $\theta$ is strictly nested, then $c_{1}(\theta)=0$.
Remark 24.
While 0-conflicts ($\bar{c}_{0}(\theta)>0$) can be defined in relation to the refinement order that gives rise to the partition lattice, the partition lattice cannot be used to detect higher-order cluster inconsistencies (1-conflicts), which can be captured and quantified instead by $\mathrm{HF}_{1}$ and the average measure $\bar{c}_{1}(\theta)$.
In Figure 2, we provide a summary of our theoretical results and their relationships.
Figure 2: Summary of key theoretical results and their relationships indicated by arrows. Double-headed arrows represent equivalences (iff), whereas single-headed arrows represent implications (if).
4.2 MCbiF as a Higher-Order Sankey Diagram
Recall the definition of the Sankey diagram of the sequence of partitions introduced in Section 3.1, and its associated representation in terms of an $M$-layered graph with vertices $V_{m}$ at each layer representing the clusters of $\theta(t_{m})$, see Eq. (1). Let us define the disjoint union $A(\ell,m):=V_{\ell}\uplus...\uplus V_{m}$, $1\leq\ell\leq m$, which assigns an index to each cluster in $\theta(t)$ for $t\in[t_{\ell},t_{m}]$. Furthermore, recall that $\theta(t)_{i}$ denotes the $i$-th cluster $C_{i}$ of $\theta(t)$. The nerve-based MCbiF can then be defined as follows.
Definition 25 (Nerve-based MCbiF).
Let $s\in[t_{\ell},t_{\ell+1}),\;\ell=1,...,M-1,$ and $t\in[t_{m},t_{m+1}),\;m=\ell,...,M-1$ or $t\geq t_{m}$ for $m=M$. We define the nerve-based MCbiF as
$\tilde{\mathcal{M}}:=(\tilde{K}^{s,t})_{t_{1}\leq s\leq t},\quad\text{where}\quad\tilde{K}^{s,t}:=\{\sigma\subseteq A(\ell,m):\bigcap_{(n,i)\in\sigma}\theta(t_{n})_{i}\neq\emptyset\}.$ (8)
The nerve-based MCbiF $\tilde{\mathcal{M}}$ is a 1-critical bifiltration with simplices representing clusters and their intersections, in contrast to the original MCbiF $\mathcal{M}$ (Eq (5)) in which the simplices represent elements in $X$ and their equivalence relations. Despite these different perspectives, Proposition 40 in the appendix shows that $\tilde{\mathcal{M}}$ and $\mathcal{M}$ lead to the same MPH and can be considered as equivalent. The proof of this equivalence follows from an extension of results by Schindler and Barahona [2025].
However, the dimensionality of $\mathcal{M}$ and $\tilde{\mathcal{M}}$ can differ, as shown in the following proposition.
Proposition 26.
  1. (i)
    $\dim K^{s,t}=\max_{s\leq r\leq t}\max_{c\in\theta(r)}|C|-1$, $\forall t_{1}\leq s\leq t$.
  2. (ii)
    $\dim\tilde{K}^{t_{m},t_{m+n}}=n,\;\forall 1\leq m\leq M,0\leq n\leq M-m$.
See proof in Appendix A.2.2. The nerve-based MCbiF is therefore computationally advantageous when $M<\max_{t_{1}\leq t}\max_{C\in\theta(t)}|C|-1$, making it preferable in applications where the clusters are large.
The nerve-based MCbiF $\tilde{\mathcal{M}}$ can be interpreted as a higher-order extension of the Sankey diagram $S(\theta)$ (Eq. 1) in the sense that $\tilde{\mathcal{M}}$ not only records pairwise intersections between clusters in consecutive partitions of $\theta$, like $S(\theta)$), but also takes into account higher-order intersections between clusters in sub-sequences of $\theta$. More formally, we can state the following proposition that relates $S(\theta)$ to the nerve-based MCbiF.
Proposition 27.
The Sankey diagram graph $S(\theta)$ is a strict 1-dimensional subcomplex of $\tilde{K}:=\tilde{K}^{t_{1},t_{M}}$. In particular, $V_{m}=K^{t_{m},t_{m}}$ and $E_{m}=K^{t_{m},t_{m+1}}$, $\forall m=1,\dots,M-1$. Hence, we can retrieve $S(\theta)$ from the zigzag filtration
$\dots\hookleftarrow\tilde{K}^{t_{m},t_{m}}\hookrightarrow\tilde{K}^{t_{m},t_{m+1}}\hookleftarrow\tilde{K}^{t_{m+1},t_{m+1}}\hookrightarrow\dots,$ (9)
which is a subfiltration of the nerve-based MCbiF.
A proof can be found in Appendix A.2.2. For details on zigzag persistence, see Carlsson and de Silva [2010] and Appendix D.
Furthermore, the 0- and 1-conflicts that can arise in a single layer $E_{m}$ of the Sankey diagram can be fully characterised as follows.
Proposition 28.
  1. (i)
    There is a 0-conflict in $t_{m},t_{m+1}]$ iff $\exists u\in V_{m}$ and $v\in V_{m+1}$ with $\mathrm{deg}(u)\geq 2$ and $\mathrm{deg}(v)\geq 2$, where $\mathrm{deg}$ denotes the node degree in the bipartite graph $(V_{m}\uplus V_{m+1},E_{m})$ associated with the Sankey diagram.
  2. (ii)
    There is a triangle 0-conflict in $t_{m},t_{m+1}]$ iff there is a path of length at least 3 in $E_{m}$.
  3. (iii)
    There is 1-conflict in $t_{m},t_{m+1}]$ iff there is an (even) cycle in $E_{m}$.
See Appendix A.2.2 for the proof and Fig. 3 for an illustration.
Figure 3: Relationship between different types of conflicts and the crossings in a single-layer Sankey diagram.
Importantly, a cycle in $E_{m}$ leads to a crossing in $E_{m}$ that cannot be undone, see Fig. 3d for an illustration. Hence, Proposition 28 (iii) implies that the sum of the elements of the superdiagonal of $\mathrm{HF}_{1}$ provides a lower bound for the minimal crossing number of the Sankey diagram, $\overline{\kappa}_{\theta}$ (defined in Eq. 2). We state this as the following corollary.
Corollary 29.
$\sum_{m=1}^{M-1}\mathrm{HF}_{1}(t_{m},t_{m+1})\leq\overline{\kappa}_{\theta}$.
Remark 30.
Note that 1-conflicts that arise across multiple partitions in the sequence (i.e., across multiple layers) do not necessarily lead to crossings. See Fig. 1, where the chosen ordering of the elements does not lead to a crossing in the layout of the Sankey diagram despite the presence of a 1-conflict. However, we hypothesise that the full $\mathrm{HF}_{0}$ and $\mathrm{HF}_{1}$ feature maps capture more complicated crossings that arise in the Sankey layout across many layers. This insight is exploited in our computational tasks below.
5 Mathematical Links of MCbiF to Other Methods
We now present some mathematical connections of MCbiF to the related methods introduced in Section 2.
Ultrametrics.
Given a sequence of partitions $\theta$ with $\theta(t_{1}=0)=\hat{0}$ and $\theta(t_{M})=\hat{1}$, let us define the matrix of first-merge times conditioned on the starting scale $s$:
$D_{\theta,s}(x_{i},x_{j}):=\min\{t\geq s\;|\;\exists\;C\in\theta(t):x_{i},x_{j}\in C\}.$ (10)
Clearly, when $s=0$, this recovers the standard matrix of first-merge times $D_{\theta}:=D_{\theta,0}$ discussed in Section 2. If $\theta$ is hierarchical, i.e., an agglomerative dendrogram, then $D_{\theta}$ is an ultrametric, i.e., it fulfils the strong triangle inequality: $D_{\theta}(x,z)\leq\max\left(D_{\theta}(x,y),D_{\theta}(x,z)\right)$ $\forall x,y,z\in X$.
Corollary 16 states that the number of branches in the agglomerative dendrogram at level $t$, which is given by $|\theta(t)|$, is equal to $\mathrm{HF}_{0}(s,t)$ for any $s\leq t$. Hence, $\mathrm{HF}_{0}(s,t)$ contains the same information as the ultrametric in the hierarchical case, see also Schindler and Barahona [2025] and Proposition 32 below.
If, on the other hand, $\theta$ is non-hierarchical, triangle 0-conflicts can lead to violations of the (strong) triangle inequality:
Proposition 31.
The triplet $x,y,z\in X$ leads to a triangle 0-conflict in $[s,t]$ iff $x,y,z$ violate the strong triangle inequality for $D_{\theta,s}$, i.e., $D_{\theta,s}(x,z)>\max(D_{\theta,s}(x,y),D_{\theta,s}(y,z))$.
See Appendix A.3 for a proof and Fig. 8a for an illustration. Proposition 31 shows that $\bar{c}_{0}(\theta)$ measures how much the ultrametric property of $D_{\theta}$ is violated.
Recall that $D_{\theta,s}$ is a dissimilarity measure that can be used to define a filtration [Chazal et al., 2014]. Next, we show that the 0-dimensional MPH of MCbiF corresponds to the 0-dimensional MPH of a Rips-based bifiltration constructed from $D_{\theta,s}$.
Proposition 32.
Let us define the Merge-Rips bifiltration $\mathcal{L}$ based on $D_{\theta,s}$ as
$\mathcal{L}=(L^{s,t})_{t_{1}\leq s\leq t}\quad\text{where}\quad L^{s,t}=\{\sigma\subset X\;\;|\;\;\forall x,y\in\sigma:\;\;D_{\theta,s}(x,y)\leq t\}.$ (11)
Then the 0-dimensional MPH of the Merge-Rips bifiltration, $\mathcal{L}$, and of the MCbiF, $\mathcal{M}$, are equivalent, but the 1-dimensional MPH of $\mathcal{L}$ and $\mathcal{M}$ are generally not equivalent. Furthermore, if $\theta$ is strictly hierarchical, then $\mathcal{L}$ has a trivial 1-dimensional MPH.
A proof is presented in Appendix A.3 and follows from an argument in Schindler and Barahona [2025]. In the hierarchical case, the 1-dimensional MPH of the Merge-Rips bifiltration is trivial because $D_{\theta}$ fulfils the strong triangle inequality, and is thus equivalent to the MCbiF, whose 1-dimensional MPH is also trivial in the hierarchical case, see Proposition 19.
Conditional Entropy.
The conditional entropy (CE) is only defined for pairs of partitions. CE is defined as the expected Shannon information of the conditional probability of a partition $\theta(t)=\{C_{1},\dots,C_{n}\}$ given $\theta(s)=\{C_{1}^{\prime},\dots,C_{m}^{\prime}\}$:
$P_{t|s}[i|j]=\frac{|C_{i}\cap C_{j}^{\prime}|}{|C_{j}^{\prime}|}.$
For the special case of $M=2$ (i.e., only two partitions in the sequence $\theta$), it can be shown that $\mathrm{HF}_{0}(t_{1},t_{2})$ follows directly from the spectral properties of the matrix $P_{t_{2}|t_{1}}P_{t_{2}|t_{1}}^{T}$ interpreted as an undirected graph, as shown by the following proposition.
Proposition 33.
$\mathrm{HF}_{0}(t_{1},t_{2})=\dim(\ker L)$, where $L=\mathrm{diag}(P_{t_{2}|t_{1}}\boldsymbol{1})-P_{t_{2}|t_{1}}P_{t_{2}|t_{1}}^{T}$ is a weighted Laplacian.
The proof can be found in Appendix A.3.
Remark 34.
Note that $P_{t|s}$ only encodes the pairwise relationship between clusters, and does not capture higher-order cluster inconsistencies. In particular, CE cannot detect 1-conflicts arising across more than two scales, as seen in Example 44 in Appendix B.
6 Experiments
6.1 Regression Task: Minimal Crossing Number of Sankey Layout
Task
In our first experiment, we consider a task of relevance in computer graphics and data visualisation: the minimisation of the crossing number of Sankey diagram layouts [Zarate et al., 2018, Li et al., 2025]. This minimisation is NP-complete, and here we use our MCbiF topological feature maps to predict the minimal crossing number $\overline{\kappa}_{\theta}$ of the Sankey diagram $S(\theta)$ of a given sequence of partitions $\theta$ (see Section 3.1).
Data
We test our measures on synthetic datasets generated by sampling randomly from the space of coarse-graining sequences of partitions.
Definition 35 (Space of coarse-graining sequences of partitions).
The space of coarse-graining sequences of partitions, denoted $\Pi_{N}^{M}$, is defined as the set of coarse-graining sequences $\theta:[0,\infty)\rightarrow\Pi_{X}$ with $|X|=N$ and $M$ change points $t_{m}=0,\ldots,M-1$, such that $|\theta(s)|\geq|\theta(t)|,\,\forall s\leq t$, which start with the partition of singletons $\theta(t_{1}=0)=\hat{0}$ and end with the full set $\theta(t_{M}=M-1)=\hat{1}$.
Setting $M=20$, we generate two datasets of 20,000 random samples $\theta\in\Pi_{N}^{M}$ for $N=5$ and $N=10$. For each of the generated $\theta$, we compute three feature maps: the information-based pairwise conditional entropy matrix $\mathrm{CE}$ [Meilă, 2003] (see Eq. 15) and our MCbiF Hilbert functions ($\mathrm{HF}_{0}$ and $\mathrm{HF}_{1}$). In addition, as a baseline feature map, we also consider the (non-unique) raw label encoding of $\theta$ given by the $N\times M$ matrix whose $m$-th column contains the labels of the clusters in $\theta(t_{m})$ assigned to the elements in $X$. As our prediction target, we take $y=\overline{\kappa}_{\theta}$ (Eq. 3), the minimal crossing number of the layout of the Sankey diagram, which we computed with the OmicsSankey algorithm [Li et al., 2025]. See Section 3.1 for details. We expect that predicting $y$ will be harder for $N=10$ because the increased complexity of $\Pi_{M}^{N}$ allows for more complicated crossings in the Sankey diagram.
Table 1: Regression task. Test $R^{2}$ score of LR, CNN and MLP models trained on different features for $N=5$ and $N=10$. See Appendix F.1 for train $R^{2}$ scores.
NN Method Raw label encoding $\mathrm{HF}_{0}$ $\mathrm{HF}_{1}$ $\mathrm{HF}_{0}$ & $\mathrm{HF}_{1}$ $\mathrm{CE}$
5 LR 0.001 0.147 0.486 0.539 0.392
CNN -0.006 0.155 0.504 0.544 0.492
MLP -0.002 0.150 0.491 0.541 0.409
10 LR -0.012 0.214 0.448 0.516 0.457
CNN 0.000 0.211 0.448 0.507 0.454
MLP 0.000 0.212 0.450 0.514 0.458
Results
As a preliminary assessment, we first compute the Pearson correlation, $r$, between the crossing number $y$ and the three measures under investigation: the information-theoretical measure $\mathrm{CI}$ (i.e., the average $\mathrm{CE}$, see Eq. (16)) and the MCbiF topological average measures $\bar{c}_{0}$ and $\bar{c}_{1}$. The correlation between $\mathrm{CI}$ and $y$ is low ($r=0.20$ for $N=5$ and $r=0.06$ for $N=10$), higher for $\bar{c}_{0}$, and highest for $\bar{c}_{1}$ ($r=0.47$ for $N=5,10$) (see Fig. 11 in Appendix F.1). This is consistent with our theoretical results in Section 4.2, which show the relation between the crossing number and $\mathrm{HF}_{1}$ (see, e.g., Corollary 29).
We then proceed to the regression task of predicting $\overline{\kappa}_{\theta}$. We split each dataset into training (64%), validation (16%) and test (20%). For each feature map (or their combinations), we train three different models: linear regression (LR), multilayer perceptron (MLP), and convolutional neural network (CNN).We use the mean-squared error (MSE) as our loss function and employ the validation set for hyperparameter tuning. See Appendix F.1 for details of all the models. We then evaluate the model performance on the unseen test data using the coefficient of determination ($R^{2}$).
We find that the raw label encoding of $\theta$ does not improve upon the mean prediction ($R^{2}\approx 0$) and that MCbiF feature maps outperform the information-based feature map (Table 1). In particular, the combined $\mathrm{HF}_{0}$ and $\mathrm{HF}_{1}$ features lead to a significantly better model performance than $\mathrm{CE}$ ($p<0.0001$, $t$-test on the residuals). Furthermore, $\mathrm{HF}_{0}$ and $\mathrm{HF}_{1}$ yield $R^{2}=0.544$ for $N=5$ and $R^{2}=0.516$ for $N=10$ whereas $\mathrm{CE}$ only achieves $R^{2}=0.492$ and $R^{2}=0.458$, respectively. The strong performance of the simple LR model demonstrates the interpretability of the MCbiF features, important for explainable AI (XAI) [Adadi and Berrada, 2018].
6.2 Classification Task: Non-Order-Preserving Sequences of Partitions
Task
In our second experiment, we classify whether a sequence of partitions is order-preserving or not, i.e., whether a sequence $\theta$ is compatible with a total ordering on the set $X$. This task is of relevance in several areas, from the study of preference relations in utility theory in social sciences [Roberts, 2009] to the analysis of weak orderings and partition refinement algorithms in computer science [Habib et al., 1999].
Definition 36 (Order-preserving sequence of partitions).
When a partition $\theta(t_{m})$ is equipped with a total order $_{m}$ on the clusters it is called an ordered partition. [^2] The ranking $\tau_{m}:V_{m}\rightarrow\{1,\dots,|V_{m}|\}$ of the vertices $V_{m}$ in the Sankey diagram $S(\theta)$ is one example of a total order $_{m}$ on the clusters, see Section 3.1. Such a partition induces a total preorder $\lesssim_{m}$ on $X$ [Stanley, 2011], i.e., if $x]_{t}<_{m}[y]_{t}$ then $x\lesssim_{m}y$. We call the sequence of partitions $\theta$ order-preserving if there exist total orders $(<_{1},\dots,<_{M})$ such that the total preorders $(\lesssim_{1},\dots,\lesssim_{M})$ are compatible across the sequence, i.e., $\forall\ell,m$ we have $x\lesssim_{\ell}y$ iff $x\lesssim_{m}y$, $\forall x,y\in X$.
Figure 4: Difference between order-preserving ($y=0$) and non-order-preserving ($y=1$) sequences (**** indicates $p<0.0001$, Mann-Whitney U test).
According to this definition, a sequence $\theta$ is non-order-preserving if there is no total order on $X$ that is consistent with all the total preorders induced by the partitions $\theta(t)$.
Data
We carry out this classification task on synthetic data for which we have a ground truth. From the space of coarse-graining sequences of partitions $\Pi_{N}^{M}$, introduced in Definition 35, we generate a balanced dataset of 3,700 partitions $\theta\in\Pi_{500}^{30}$, half of which are order-preserving ($y=0$) and the other half are non-order-preserving ($y=1$). The loss of order-preservation is induced by introducing random swaps in the node labels across layers. See Appendix F.2 for details. For each of the generated $\theta$ we compute $\mathrm{CE}$, $\mathrm{HF}_{0}$ and $\mathrm{HF}_{1}$ using the computationally advantageous nerve-based MCbiF. We choose $N=500$ and $M=30$ to demonstrate the scalability of our method.
Table 2: Classification task. Test accuracy of logistic regression trained on different features.
Raw label
encoding $\mathrm{HF}_{0}$ $\mathrm{HF}_{1}$ $\mathrm{CE}$
0.53 0.56 0.97 0.50
Results
Firstly, whereas we find no significant difference between the information-theoretical $\mathrm{CI}$ of order-preserving ($y=0$) and non-order-preserving ($y=1$) sequences, we observe a statistically significant increase of $\bar{c}_{0}$ and $\bar{c}_{1}$ for order-preserving sequences (Fig. 4).
For the classification task, we split our data into training (80%) and test (20%). For each feature map, we then train a logistic regression on the training split, and evaluate the accuracy on the test split, see Appendix F.2. We find that $\mathrm{HF}_{1}$ predicts the label $y=\{0,1\}$ encoding the (lack of) order preservation with high accuracy (0.95). In contrast, $\mathrm{CE}$ and the raw label encoding of $\theta$ cannot improve on a random classifier (Table 2). Our results thus demonstrate the high sensitivity of MCbiF to order-preservation in $\theta$ because non-order-preserving sequences induce 1-conflicts that we capture with $\mathrm{HF}_{1}$.
6.3 Application to Real-World Temporal Data
In our final experiment, we apply MCbiF to temporal sequences of partitions computed from real-world contact data of free-ranging house mice that capture the changes in the social network structure of the rodents over time [Bovet et al., 2022].
Data
Each partition $\theta_{\tau}(t)$ describes mice social groupings for $N=281$ individual mice at weeks $t\in[1,\dots,9]$, i.e., the nine weeks in the study period (28 February-1 May 2017). Hence, each sequence captures the fine-graining of social groups over the transition from winter to spring. Each partition sequence is computed at temporal resolution $\tau>0$, where the parameter $\tau$ modulates how fine the temporal community structure is (Fig. 13). See Bovet et al. [2022] for details.
We use MCbiF to compare the temporal sequences $\theta_{\tau_{i}}$ for nine parameters $\tau_{i}$, $i=1,\dots,9$, as provided in Bovet et al. [2022]. See Appendix F.3 for details. For each of the nine partition sequences $\theta_{\tau_{i}}$, we compute $\mathrm{HF}_{0}$ and $\mathrm{HF}_{1}$ using the computationally advantageous nerve-based MCbiF, which induces a 50-fold reduction in computation time due to a much lower number of simplices (260 simplices for the nerve-based MCbiF instead of 116,700 for the original MCbiF).
Figure 5: (a) Analysis of non-hierarchical sequences of partitions $\theta_{\tau_{i}}$ compiled from the temporal social interactions of a mice population over a period of 9 weeks. Each $\theta_{\tau_{i}}$ is formed by a sequence of social groupings $\theta_{\tau_{i}}(t)$ over week $t$. Different sequences of partitions were computed as a function of the parameter $\tau_{i}$. We display Sankey diagrams and MCbiF feature maps for $\theta_{\tau_{i}}$ at three parameters $\tau_{i}$ ($i=2,4,8$) identified as robust in the original work by Bovet et al. [2022]. These three sequences $\theta_{\tau_{i}}$ exhibit different types of non-hierarchy, as shown by our topological feature maps and our measures of average 0-conflict ($\bar{c}_{0}$) and average 1-conflict ($\bar{c}_{1}$). (b) The $\theta_{\tau_{i}}$ ($i=2,4,8$) found in Bovet et al. [2022] as robust behaviours correspond to distinct topological characteristics of the sequences of partitions, as captured by the block structure in the distance between MCbiF Hilbert functions.
Results
Bovet et al. [2022] identified that the temporal resolutions $\tau_{2}=1\,\mathrm{s}$, $\tau_{4}=60\,\mathrm{s}$ and $\tau_{8}=24\,\mathrm{h}$ lead to robust sequences of partitions. Using the Hilbert distance, i.e., the $L_{2}$-norm on the 0- and 1-dimensional MCbiF Hilbert functions, we find these temporal resolutions to be representative for three distinct temporal regimes characterised by different degrees of non-hierarchy, as measured by $\bar{c}_{0}$ and $\bar{c}_{1}$ (Fig. 5). In particular, high $\bar{c}_{0}$ indicates that mice tend to split off groups over time, and high $\bar{c}_{1}$ indicates that mice meet in overlapping subgroups but never jointly in one nest box. Note that $\theta_{\tau_{2}}$ has a strong non-hierarchical structure because the large-scale mice social clusters get disrupted in the transition to spring. In contrast, $\theta_{\tau_{8}}$ is more hierarchical as it captures the underlying stable social groups revealed by the higher temporal resolution. However, $\theta_{\tau_{4}}$ has the strongest hierarchy as indicated by a lower $\bar{c}_{0}$ and an absence of 1-conflicts ($\bar{c}_{1}=0$) and thus corresponds to a sweet spot in hierarchical organisation between the low and high temporal resolutions. Finally, the Hilbert distances also capture an increased time reversibility in the sequence $\theta_{\tau_{8}}$ due to the increased stability of social groupings over time, see Fig. 14 in Appendix F.3.
7 Conclusion
We have introduced the MCbiF, a novel bifiltration that encodes the cluster intersection patterns in a multiscale, non-hierarchical sequence of partitions, $\theta$. Its stable Hilbert functions $\mathrm{HF}_{k}$ quantify the topological autocorrelation of $\theta$ and measure non-hierarchy in two complementary ways: the Hilbert function at dimension $k=0$ captures the absence of a maximum with respect to the refinement order (0-conflicts), whereas the Hilbert function at dimension $k=1$ captures the emergence of higher-order cluster inconsistencies (1-conflicts). This is summarised by the measures of average 0-conflict $\bar{c}_{0}(\theta)$ and average 1-conflict $\bar{c}_{1}(\theta)$, which are global, history-dependent and sensitive to the ordering of the partitions in $\theta$. The MCbiF extends the 1-parameter MCF defined by Schindler and Barahona [2025] to a 2-parameter filtration, leading to richer algebraic invariants that describe the full topological information in $\theta$. We remark that the MCbiF is independent of the chosen clustering algorithm and can be applied to any (non-hierarchical) sequence of partitions $\theta$.
We demonstrate with numerical experiments that the MCbiF Hilbert functions provide topological feature maps that can be used for downstream machine learning tasks, and are shown to outperform information-based features on regression and classification tasks on non-hierarchical sequences of partitions. Moreover, the grounding of MCbiF features in algebraic topology enhances interpretability, a crucial attribute for XAI and applications to real-world data.
Limitations and future work
Our analysis of the MCbiF MPH is restricted to dimensions 0 and 1 due to current limitations of the RIVET software [Lesnick and Wright, 2015] used in our numerical experiments. Analysing topological autocorrelation for higher dimensions would allow us to capture more complex higher-order cluster inconsistencies and could be the object of future research.
Furthermore, we focused here on Hilbert functions as our topological invariants because of their computational efficiency and analytical simplicity, which facilitates our theoretical analysis. In future work, we plan to use richer feature maps by exploiting the block decomposition of the MCbiF persistence module, which leads to barcodes [Bjerkevik, 2021], or by using multiparameter persistence landscapes [Vipond, 2020].
Another future direction is to use MCbiF to evaluate the consistency of assignments in consensus clustering [Strehl and Ghosh, 2002, Vega-Pons and Ruiz-Shulcloper, 2011]. Indeed, it can be shown that the values of the Hilbert function $\mathrm{HF}_{k}(s,t)$ that are further away from the diagonal ($s=t$) are more robust to permutations of the ordering of partitions in $\theta$ (see Proposition 41 in Appendix A.4), and, in particular, $\mathrm{HF}_{k}(t_{1},t_{M})$ only depends on the set of distinct partitions in the sequence $\theta([t_{1},\infty))$ and is independent to any permutation in their order. Hence, in future work, $\mathrm{HF}_{k}(t_{1},t_{M})$ could be used as an overall measure of consistency in $\theta$ in the context of consensus clustering.
Finally, we plan to analyse minimal cycle representatives of the MPH [Li et al., 2021] to localise 1-conflicts in the sequence of partitions, which is of interest to compute conflict-resolving partitions in consensus clustering, or to identify inconsistent assignments in temporal clustering [Liechti and Bonhoeffer, 2020].
Reproducibility Statement
Detailed proofs of all theoretical results can be found in Appendix A. Extensive documentation of our experiments is presented in Appendix F. The dataset studied in Section 6.3 is publicly available at: https://dataverse.harvard.edu/file.xhtml?fileId=5657692. Python code for the MCbiF method is publicly available at: https://github.com/barahona-research-group/MCF
Acknowledgments
JS acknowledges support from the EPSRC (PhD studentship through the Department of Mathematics at Imperial College London). MB acknowledges support from EPSRC grant EP/N014529/1 supporting the EPSRC Centre for Mathematics of Precision Healthcare. We thank Kevin Michalewicz, Asem Alaa and Christian Schindler for valuable discussions on the computational aspects of this project. We thank Alex Bovet for discussions about the temporal dataset studied in this paper. We also thank Arne Wolf for discussions about the stability of multiparameter persistence modules. This work benefited from discussions at the London-Oxford TDA seminar in November 2024.
Appendices
Appendix A Proofs of Theoretical Results
A.1 Proofs for Section 1
We first state a simple fact about coarse-graining sequences of partitions.
Remark 37.
Let $\theta(t)_{i}$ denote the $i$-th cluster $C_{i}$ of $\theta(t)$. It is a simple fact that $\theta$ is coarse-graining iff the mean cluster size is non-decreasing, i.e., $\frac{1}{|\theta(s)|}\sum_{i}^{|\theta(s)|}|\theta(s)_{i}|\leq\frac{1}{|\theta(t)|}\sum_{j}^{|\theta(t)|}|\theta(t)_{j}|$ for $s\leq t$. The proof follows directly from the fact that $\sum_{i}^{|\theta(s)|}|\theta(s)_{i}|=\sum_{i}^{|\theta(t)|}|\theta(t)_{i}|=N$.
A.2 Proofs for Section 4
We provide a proof for the multi-criticality of the MCbiF filtration stated in Proposition 4.
Proof of Proposition 4.
The MCbiF is indeed a bifiltration because $K^{s,t}\subseteq K^{s^{\prime},t^{\prime}}$ if $s\geq s^{\prime}$ and $t\leq t^{\prime}$. See Fig. 6 for the triangular diagram of the MCbiF filtration, where arrows indicate inclusion maps. The MCbiF is uniquely defined by its values on the finite grid $t_{1},\dots,t_{M}]\times[t_{1},\dots,t_{M}]$ because $\theta$ has change points $t_{1}<\dots<t_{M}$. It is a multi-critical bifiltration because for $x\in X$ we have $x]\in K^{s,t}$ for all $s,t\in[t_{1},\infty)^{\mathrm{op}}\times[t_{1},\infty)$. In particular, $x\in K^{t_{1},t_{1}}$ and $x\in K^{t_{1}+\delta,t_{1}+\delta}$ for $\delta>0$ but $(t_{1},t_{1})$ and $(t_{1}+\delta,t_{1}+\delta)$ are incomparable in the poset $t_{1},\infty)^{\mathrm{op}}\times[t_{1},\infty)$. ∎
$\dots$$\dots$$\dots$$K^{t,t}$$K^{t,t^{\prime}}$$K^{t,t^{\prime\prime}}$$\dots$$K^{t^{\prime},t^{\prime}}$$K^{t^{\prime},t^{\prime\prime}}$$\dots$$K^{t^{\prime\prime},t^{\prime\prime}}$$\dots$}
Figure 6: Triangular commutative diagram of the MCbiF for $t_{1}\leq t\leq t^{\prime}\leq t^{\prime\prime}$. The arrows indicate inclusion maps between simplicial complexes.
Next, we provide formal definitions for algebraic properties of persistence modules, see Botnan and Lesnick [2023] for details.
Definition 38.
For partially ordered sets $P_{1},P_{2}$, we call an interval $I\subseteq P_{1}\times P_{2}$ a block if it can be written as one of the following types:
  1. 1.
    Birth quadrant: $I=S_{1}\times S_{2}$ for downsets $S_{1}\subseteq P_{1}$ and $S_{2}\subseteq P_{2}$.
  2. 2.
    Death quadrant: $I=S_{1}\times S_{2}$ for upsets $S_{1}\subseteq P_{1}$ and $S_{2}\subseteq P_{2}$.
  3. 3.
    Vertical band: $I=S_{1}\times P_{2}$ for an interval $S_{1}\subseteq P_{1}$.
  4. 4.
    Horizontal band: $I=P_{1}\times S_{2}$ for an interval $S_{2}\subseteq P_{2}$.
Definition 39.
Let $\mathbf{Vect}$ denote the category of $\boldsymbol{k}$-vector spaces for a fixed field $\boldsymbol{k}$. For a partially ordered set $P$, a $P$-indexed persistence module is a functor $F:P\mapsto\mathbf{Vect}$. We say that:
  1. a)
    $F$ is called pointwise finite-dimensional if $\dim(F_{\boldsymbol{a}})<\infty$ for all $\boldsymbol{a}\in P$.
  2. b)
    $F$ is called finitely presented if there exists a morphism of free modules $\phi_{1}:F_{1}\rightarrow F_{1}$ such that $\mathrm{coker}(\phi_{1})\cong F$ and $F_{0}$ and $F_{1}$ are finitely generated.
  3. c)
    $F$ is called block-decomposable if it decomposes into blocks $F\bigoplus_{B\in\mathcal{B}(F)}\boldsymbol{k}_{B}$ where $\mathcal{B}(F)$ is a multiset of blocks that depends on $F$.
We can now provide the proof for Proposition 6, which shows that the MCbiF persistence module (see Fig. 7) is pointwise finite-dimensional, finitely presented and block-decomposable. The proof relies on the equivalent nerve-based construction of the MCbiF (see Proposition 40), and the exactness of the persistence module from which block-decomposability follows [Cochoy and Oudot, 2020].
Proof of Proposition 6.
The MCbiF module is pointwise finite-dimensional because the homology groups of finite simplicial complexes are finite. As the MCbiF is defined uniquely by its values on a finite grid (Proposition 4), its persistence module consists of finitely many vector spaces and finitely many linear maps between them, hence it is finitely presented.
To prove block-decomposability, we use the nerve-based MCbiF $(\tilde{K}^{s,t})_{t_{1}\leq s\leq t}$, which leads to the same persistence module, see Proposition 40. As the module is uniquely defined by its values on a finite grid, we can use Theorem 9.6 by Cochoy and Oudot [2020] that implies block-decomposability if the persistence module is exact. Hence, it suffices to show that for all $t_{1}\leq t\leq t^{\prime}\leq t^{\prime\prime}\leq t^{\prime\prime\prime}$ the diagram
$H_{k}(\tilde{K}^{t,t^{\prime\prime}})$$H_{k}(\tilde{K}^{t,t^{\prime\prime\prime}})$$H_{k}(\tilde{K}^{t^{\prime},t^{\prime\prime}})$$H_{k}(\tilde{K}^{t^{\prime},t^{\prime\prime\prime}})$}
induces an exact sequence:
$H_{k}(\tilde{K}^{t^{\prime},t^{\prime\prime}})\rightarrow H_{k}(\tilde{K}^{t,t^{\prime\prime}})\oplus H_{k}(\tilde{K}^{t^{\prime},t^{\prime\prime\prime}})\rightarrow H_{k}(\tilde{K}^{t,t^{\prime\prime\prime}})$ (12)
By construction of the MCbiF, $\tilde{K}^{t,t^{\prime\prime\prime}}=\tilde{K}^{t,t^{\prime\prime}}\cup\tilde{K}^{t^{\prime},t^{\prime\prime\prime}}$. Furthermore, $\tilde{K}^{t,t^{\prime\prime}}=\tilde{K}^{t,t^{\prime}}\cup\tilde{K}^{t^{\prime},t^{\prime\prime}}$ and $\tilde{K}^{t^{\prime},t^{\prime\prime\prime}}=\tilde{K}^{t^{\prime},t^{\prime\prime}}\cup\tilde{K}^{t^{\prime\prime},t^{\prime\prime\prime}}$. Without loss of generality, $_{k}, t=tt^{\prime}=t_{\ell}$, $^{\prime\prime}=t_{m}, t′′′=t$$^{\prime\prime\prime}=t_{n}$ for change points $t_{k}<t_{\ell}<t_{m}<t_{n}$ of $\theta$ such that $A(k,\ell)\cap A(m,n)=\emptyset$. Hence, $\tilde{K}^{t,t^{\prime}}\cap\tilde{K}^{t^{\prime\prime},t^{\prime\prime\prime}}=\emptyset$ and $\tilde{K}^{t^{\prime},t^{\prime\prime}}=\tilde{K}^{t,t^{\prime\prime}}\cap\tilde{K}^{t^{\prime},t^{\prime\prime\prime}}$. This means that Eq. (12) is a Mayer-Vietoris sequence for all $k\geq 0$, implying exactness [Hatcher, 2002, p. 149] and proving the block decomposability [Cochoy and Oudot, 2020, Theorem 9.6]. ∎
$\dots$$\dots$$\dots$$H_{k}(K^{t,t})$$H_{k}(K^{t,t^{\prime}})$$H_{k}(K^{t,t^{\prime\prime}})$$\dots$$H_{k}(K^{t^{\prime},t^{\prime}})$$H_{k}(K^{t^{\prime},t^{\prime\prime}})$$\dots$$H_{k}(K^{t^{\prime\prime},t^{\prime\prime}})$$\dots$}
Figure 7: Multiparameter persistence module of the MCbiF for $t_{1}\leq t\leq t^{\prime}\leq t^{\prime\prime}$. The arrows indicate linear maps between vector spaces.
A.2.1 Proofs for Section 4.1
We continue with the proof of Proposition 12 that relates 0-conflicts to hierarchy and triangle 0-conflicts to nestedness.
Proof of Proposition 12.
(ii) If $\theta$ has a 0-conflict then $\exists r_{1},r_{2}\in[s,t]$ such that $\theta(r_{1})\nleq\theta(r_{2})$ and $\theta(r_{1})\ngeq\theta(r_{2})$, otherwise $\theta([s,t])$ would have a maximum. Hence, $\theta$ is not hierarchical in $[s,t]$.
(iii) Let us first assume that $\theta$ is coarse-graining, i.e., $|\theta(t)|\leq|\theta(r)|$ for all $r\in[s,t]$. We show that no 0-conflict in $[s,t]$ implies that $\theta$ is hierarchical in $[s,t]$. Let $r_{1},r_{2}\in[s,t]$ with $r_{1}\leq r_{2}$, then the subposet $\theta([r_{1},r_{2}])$ has a maximum because of the absence of a 0-conflict, and the maximum is given by $\theta(r_{2})$ due to coarse-graining. Hence, $\theta(r_{1})\leq\theta(r_{2})$. As $r_{1},r_{2}$ were chosen arbitrarily, this implies that $\theta$ is hierarchical in $[s,t]$. The argument is analogous for the case that $\theta$ is fine-graining.
(iv) Let $x,y,z\in X$ be in a triangle 0-conflict. In particular, $x\sim_{r_{1}}y\sim_{r_{2}}z$ with $x\neq y$, $x\neq z$ and $y\neq z$. Hence, there are $C\in\theta(r_{1})$ and $C^{\prime}\in\theta(r_{2})$ such that $x,y\in C$ and $y,z\in C^{\prime}$, as well as $z\notin C$ and $x\notin C^{\prime}$. This implies $\{x\}\in C\setminus C^{\prime}$, $\{z\}\in C^{\prime}\setminus C$ and $\{y\}\in C\cap C^{\prime}$, showing that $C$ and $C^{\prime}$ are non-nested. Hence, $\theta$ is non-nested in $[s,t]$.
(i) Moreover, $\nexists r\in[s,t]$ such that $x\sim_{r}\sim_{r}y$. In particular, $\nexists r\in[s,t]$ such that $\nexists C^{\prime\prime}\in\theta(r)$ with $C\subseteq C^{\prime\prime}$ and $C\subseteq C^{\prime\prime}$. Hence, $\nexists r\in[s,t]$ such that $\theta(r_{1})\leq\theta(r)$ and $\theta(r_{2})\leq\theta(r)$, implying that the subposet $\theta([s,t])$ has no maximum. This shows that every triangle 0-conflict is also a 0-conflict, proving statement (i). Note that the opposite is not true as illustrated by the example in Fig. 3b. ∎
We now provide a proof for Proposition 14 on properties of the 0-dimensional Hilbert function of the MCbiF.
Proof of Proposition 14.
(i) $\mathrm{HF}_{0}(s,t)$ is equal to the number of connected components of $K^{s,t}$. Let $r^{\prime}\in[s,t]$ such that $c=|\theta(r^{\prime})|=\min_{r\in[s,t]}|\theta(r)|$. We can represent $\theta(r)=\{C_{1},\dots,C_{c}\}$ and by construction $\Delta C\in K^{s,t}$ for all $C\in\theta(r)$. Hence, if two elements $x,y\in X$ are in the same cluster $C\in\theta(r)$ then $x,y]\in K^{s,t}$ and the 0-simplices $x],[y]\in K^{s,t}$ are in the same connected component. As $\theta(r)$ has $c$ mutually disjoint clusters, this means that there cannot be more than $c$ disconnected components in $K^{s,t}$ and $\mathrm{HF}_{0}(s,t)\leq c=|\theta(r^{\prime})|$. As $r^{\prime}\in[s,t]$ was chosen arbitrarily, this implies $\mathrm{HF}_{0}(s,t)\leq\min_{r\in[s,t]}|\theta(r)|$.
We prove statement (ii) by the contrapositive and show that the following two conditions are equivalent:
  1. C1:
    $\mathrm{HF}_{0}(s,t)=\min_{r\in[s,t]}|\theta(r)|$.
  2. C2:
    $\exists r\in[s,t]$ such that $\theta(r^{\prime})\leq\theta(r)$, $\forall r^{\prime}\in[s,t]$.
[…]
$\Longrightarrow$” To prove the other direction, assume that C1 is true. Then there exists $r\in[s,t]$ such that $c:=\mathrm{HF}_{0}(s,t)=|\theta(r)|$ with $|\theta(r)|=\min_{q\in[s,t]}|\theta(q)|$. In particular, the disconnected components of $K^{s,t}$ are given by the clusters of $\theta(r)$ denoted by $C_{1},\dots,C_{c}$. Let $r^{\prime}\in[s,t]$ and $C^{\prime}\in\theta(r)$. Then $\exists i\in[1,\dots,c]$ such that $C^{\prime}\subseteq C_{i}\in\theta(r)$ because otherwise the solid simplex $\Delta C^{\prime}$ would connect two solid simplices in $\{\Delta C_{1},\dots,\Delta C_{c}\}$, contradicting that they are disconnected in $K^{s,t}$. Hence, the clusters of $\theta(r^{\prime})$ are all subsets of cluster of $\theta(r)$, implying $\theta(r^{\prime})\leq\theta(r)$. As $r^{\prime}\in[s,t]$ was chosen arbitrary this shows C2.
[…]
We next prove Corollary 16 about some properties of the average 0-conflict, which follows immediately from Proposition 14.
Proof of Corollary 16.
We begin with the proof of statement (i). If $\theta$ is hierarchical in $[s,t]$ then $\theta$ is either coarse- or fine-graining. Assume first that $\theta$ is coarse-graining, then $\theta(s^{\prime})\leq\theta(t)$ for all $s^{\prime}\in[s,t]$ and together with hierarchy, this implies that $\theta(t)$ is an upper bound of the subposet $\theta([s,t])$. Hence, Proposition 14 (iii) shows that $\mathrm{HF}_{0}(s,t)=|\theta(t)|$. Moreover, $\mathrm{HF}_{0}(s,t)=\min(|\theta(s)|,|\theta(t)|)$ because coarse-graining implies $|\theta(s)|\geq|\theta(t)|$. A similar argument also shows $\mathrm{HF}_{0}(s,t)=|\theta(s)|=\min(|\theta(s)|,|\theta(t)|)$ if $\theta$ is fine-graining.
We continue with proving (ii). $\bar{c}_{0}(\theta)>0$ is equivalent to $\exists s,t\in[t_{1},t_{M}]$ such that $\mathrm{HF}_{0}(s,t)<\min_{r\in[s,t]}|\theta(r)|$, according to Definition 15. This is again equivalent to $\exists s,t\in[t_{1},t_{M}]$ such that $\theta$ has a 0-conflict in $[s,t]$, according to Proposition 14 (ii).
[…]
Next, we provide the proof of Proposition 17 about the relation between triangle 0-conflicts and the clustering coefficient.
Proof of Proposition 17.
Assume that $\mathcal{C}(K_{1}^{s,t})<1$. Then there exist $x,y,z\in X$ that form a path of length 2 but no triangle, see Newman [2018] for details on the clustering coefficient. Without loss of generality, $x,y],[y,z]\in K_{1}^{s,t}$ but $x,z]\notin K_{1}^{s,t}$. This implies $\exists r_{1},r_{2}\in[s,t]$: $x\sim_{r_{1}}y\sim_{r_{2}}z$ and $\nexists r\in[s,t]$: $x\sim_{r}z$. Hence, $x,y,z$ lead to a triangle 0-conflict. ∎
We can now prove Proposition 19 on 1-conflicts.
Proof of Proposition 19.
Statement (i) follows directly from the definition of 1-conflicts that $\mathrm{HF}_{1}(s,t)=\dim[H_{k}(K^{s,t})]\geq 1$ iff $\theta$ has a 1-conflict.
We next prove statement (ii): If $\mathrm{HF}_{1}(s,t)\geq 1$ there exists a 1-cycle $z=[x_{1},x_{2}]+\dots+[x_{n-1},x_{n}]+[x_{n},x_{1}]$ that is non-bounding, i.e., $h:=[z]\neq 0$ in $H_{1}(K^{s,t})$, see Appendix C for details. Case 1: Assume $\nexists r\in[s,t]$: $x_{1}\sim_{r}x_{2}\sim_{r}x_{3}$, then it follows immediately that $x_{1},x_{2},x_{3}$ lead to a triangle 0-conflict. Case 2: Assume $\exists r\in[s,t]$: $x_{1}\sim_{r}x_{2}\sim_{r}x_{3}$. As $z]\neq 0$ there exists a 1-cycle $\tilde{z}=[\tilde{x}_{1},\tilde{x}_{2}]+\dots++[\tilde{x}_{m-1},\tilde{x}_{m}]+[\tilde{x}_{m},\tilde{x}_{1}]\in Z_{1}(K^{s,t})$ such that $\tilde{z}$ is homologous to $z$, i.e., $\tilde{z}=z+\partial_{2}w$ for $w\in C_{2}(K^{s,t})$, and such that $\nexists r\in[s,t]$: $\tilde{x}_{1}\sim_{r}\tilde{x}_{2}\sim_{r}\tilde{x}_{3}$. In particular, $\tilde{x}_{1},\tilde{x}_{2},\tilde{x}_{3}$ lead to a triangle 0-conflict.
We finally prove statement (iii): If $\theta$ is hierarchical, then it has no 0-conflicts according to Corollary 16. Hence, $\theta$ also has no triangle 0-conflict in $[s,t]$ and so (i) implies that $\mathrm{HF}_{1}(s,t)=0$. ∎
A.2.2 Proofs for Section 4.2
Next, we provide the proof about the equivalence between MCbiF and nerve-based MCbiF.
Proposition 40.
The bifiltrations $\mathcal{M}$ and $\tilde{\mathcal{M}}$ lead to the same persistence module.
Proof.
The proof follows from Proposition 30 in Schindler and Barahona [2025], which extends directly to the 2-parameter case. ∎
Next, we prove Proposition 26 about the dimension of the nerve-based MCbiF.
Proof of Proposition 26.
Statement (i) follows directly from the definition in Eq. (5). We show statement (ii) by induction. Base case: From the definition of the nerve-based MCbiF, it follows directly that $\dim N^{t_{m},t_{m}}=0$ because the indices in $A(m,m)$ correspond to mutually exclusive clusters. Induction step: Let us assume that $\dim N^{t_{m},t_{m+n}}=n$, then there exist $C_{0},\dots,C_{n}\in\theta([t_{m},t_{m+n}])$ such that $C_{0}\cap\dots\cap C_{n}\neq\emptyset$. As the clusters in partition $\theta(t_{m+n+1})$ cover the set $X$ there exist a cluster $C\in\theta(t_{m+n+1})$ such that $C\cap C_{0}\cap\dots\cap C_{n}\neq\emptyset$. Hence, $\dim N^{t_{m},t_{m}+n}\geq n+1$. If $\dim N^{t_{m},t_{m}+n}>n+1$ there would exist a second cluster $C^{\prime}\in\theta(t_{m+n+1})$ with $C^{\prime}\cap C\cap C_{0}\cap\dots\cap C_{n}\neq\emptyset$ but $C^{\prime}\cap C\neq\emptyset$ contradicts that clusters of $\theta(t_{m+n+1})$ are mutually exclusive. Hence, $\dim N^{t_{m},t_{m}+n}=n+1$, proving statement (ii) by induction. ∎
We provide a proof for the connection between Sankey diagrams and the nerve-based MCbiF.
Proof of Proposition 27.
The Sankey diagram graph $S(\theta)=(V=V_{1}\uplus...\uplus V_{M},E=E_{1}\uplus...\uplus E_{M-1})$ is a strict 1-dimensional subcomplex of $\tilde{K}=\tilde{K}^{t_{1},t_{M}}$ because $\tilde{K}^{t_{m},t_{m}}=V_{m}\subseteq\tilde{K}$ and $\tilde{K}^{t_{m},t_{m+1}}=E_{m}\subseteq\tilde{K}$. This also shows that the zigzag filtration (9) contains exactly the same vertices (0-simplices) and edges (1-simplices) as $S(\theta)$. ∎
We next prove Proposition 28 that characterises conflicts that can arise in a single layer of the Sankey diagram.
Proof of Proposition 28.
(i) Suppose that $\theta$ has a 0-conflict in $t_{m},t_{m+1}]$. Then $\theta(t_{m})\nleq\theta(t_{m+1})$ and $\theta(t_{m})\ngeq\theta(t_{m+1})$. This means that there exists $C\in\theta(r_{1})$ such that $\exists C^{\prime},C^{\prime\prime}\in\theta(r_{2})$ with $\cap C^{\prime}\neq\emptyset, CC$$\cap C^{\prime\prime}\neq\emptyset$ and $C^{\prime}\cap C^{\prime\prime}=\emptyset$, otherwise $\theta(r_{1})\leq\theta(r_{2})$. Hence, $\theta(r_{1})\nleq\theta(r_{2})$ is equivalent to $\exists u\in V_{m}$ (the node corresponding to cluster $C$) with degree $\mathrm{deg}(u)\geq 2$ in $E_{m}$. An analogous argument shows that $\theta(r_{1})\ngeq\theta(r_{2})$ is equivalent to $\exists v\in V_{m+1}$ with $\mathrm{deg}(v)\geq 2$. This proves the statement.
(ii) Let $x,y,z\in X$ form a triangle 0-conflict for the interval $t_{m},t_{m+1}]$, i.e., $x\sim_{t_{m}}y\sim_{t_{m+1}}z$ but $x\nsim_{t_{m+1}}y\nsim_{t_{m}}z$. In particular, the elements $x,y,z$ are mutually distinct. This means there exist $C_{1},C_{2}\in\theta(t_{m})$ and $C_{1}^{\prime},C_{2}^{\prime}\in\theta(t_{m+1})$ such that $x,y\in C_{1}$, $z\in C_{2}$, $y,z\in C_{1}^{\prime}$ and $x\in C_{2}^{\prime}$. This is equivalent to $C_{1}\cap C_{2}=\emptyset$, $C_{1}^{\prime}\cap C_{2}^{\prime}=\emptyset$ and $C_{1}\cap C_{1}^{\prime}\neq\emptyset$, $C_{1}\cap C_{2}^{\prime}\neq\emptyset$, $C_{2}\cap C_{2}^{\prime}\neq\emptyset$. Let $u,u^{\prime}\in V_{m}$ correspond to $C_{1}$ and $C_{2}$, respectively, and $v,v^{\prime}\in V_{m+1}$ correspond to $C_{1}^{\prime}$ and $C_{2}^{\prime}$, respectively. Then the above is equivalent to $u^{\prime},v],[v,u],[u,v^{\prime}]\in E_{m}$, which is again equivalent to the existence of a path in $E_{m}$ that has length at least 3.
(iii) The statement follows from the fact that every cycle in $E_{m}$ is even because the graph $(V_{m}\uplus V_{m+1},E_{m})$ is bipartite and the fact that every cycle in $E_{m}=K^{t_{m},t_{m+1}}$ is non-bounding because $\dim K^{t_{m},t_{m+1}}=1$. ∎
A.3 Proofs for Section 5
We continue by proving that 0-conflicts can induce violations of the strong triangle inequality as stated in Proposition 31.
Proof of Proposition 31.
Let $x,y,z\in X$ lead to a triangle 0-conflict for the interval $[s,t]$, i.e., $\exists r_{1},r_{2}\in[s,t]$: $x\sim_{r_{1}}y\sim_{r_{2}}z$ and $\nexists r\in[s,t]$: $x\nsim_{r}y\nsim_{r}z$. This means $D_{\theta,s}(x,y)\leq r_{1}$ and $D_{\theta,s}(y,z)\leq r_{2}$. Let us define $r_{3}:=D_{\theta,s}(x_{z})\leq t_{M}$. We know that $r_{3}\neq r_{1}$ and $r_{3}\neq r_{2}$ as otherwise $x\sim_{r_{3}}y\sim_{r_{3}}z$ for $r_{3}\in[s,t]$, contradicting the lack of transitivity. Hence, without loss of generality, $r_{1}<r_{2}<r_{3}$. Then the above is equivalent to $D_{\theta,s}(x_{z})=r_{3}>r_{1}\geq\min(D_{\theta,s}(x,y),D_{\theta,s}(y,z))$, which is a violation of the strong triangle inequality. ∎
Next, we provide a proof of Proposition 32 that establishes the connection between the MPH of the MCbiF and that of the Merge-Rips bifiltration constructed from the matrix of first-merge times, $D_{\theta,s}$.
Proof of Proposition 32.
First note that $\mathcal{L}=(L^{s,t})_{t_{1}\leq s\leq t}$ is indeed a well-defined bifiltration because $L^{s,t}\subseteq L^{s^{\prime},t^{\prime}}$ if $s\geq s^{\prime}$ and $t\leq t^{\prime}$. In particular, $\mathcal{L}$ is also defined uniquely on the finite grid $P=\{(s,t)\in[t_{1},\dots,t_{M}]\times[t_{1},\dots,t_{M}]\;|\;s\leq t\}$ with partial order $(s,t)\leq(s^{\prime},t^{\prime})$ if $s\geq s^{\prime},t\leq t^{\prime}$.
The proof of the proposition then follows from a simple extension of Proposition 32 in Schindler and Barahona [2025] to the 2-parameter case. To see that the 0-dimensional MPH of $\mathcal{L}$ and $\mathcal{M}$ are equivalent, note that both bifiltrations have the same 1-skeleton. Moreover, the 1-dimensional MPH is generally not equivalent because $\mathcal{L}$ is a Rips-based bifiltration and thus 2-determined, whereas $\mathcal{M}$ is not 2-determined.
If $\theta$ is strictly hierarchical, then $D_{\theta,s}$ fulfils the strong-triangle inequality and thus the Rips-based bifiltration leads to a trivial 1-dimensional homology, see [Schindler and Barahona, 2025, Corollary 33]. Hence, the 1-dimensional MPH of $\mathcal{L}$ is trivial. ∎
Finally, we provide a brief proof for Proposition 33 linking the 0-dimensional Hilbert function of a pair of partitions and the graph Laplacian built from the conditional entropy matrix between both partitions.
Proof of Proposition 33.
Using Proposition 27, we prove the statement with the equivalent nerve-based MCbiF. Note that the graph $G:=P_{t_{2}|t_{1}}P_{t_{2}|t_{1}}^{T}$ has the same vertices and edges as the simplicial complex $\tilde{K}^{t_{1},t_{2}}$, which is 1-dimensional and thus also a graph according to Proposition 26. This shows that $\mathrm{HF}_{(}t_{1},t_{2})$ is given by the number of connected components in $G$. Furthermore, observe that $P_{t_{2}|t_{1}}^{T}\mathbf{1}=\mathbf{1}$, and that the resulting matrix $L$ is the Laplacian of the undirected graph $G$. Hence, $\dim(\ker L)$ is equal to the number of connected graph components [Chung, 1997], proving the statement. ∎
A.4 Proofs for Section 7
It follows from the construction of MCbiF that the Hilbert functions are invariant to certain swaps of partitions in $\theta$.
Proposition 41.
$\mathrm{HF}_{k}(s,t)$ is invariant to swaps of partitions in sequence $\theta$ between $s$ and $t$, for $t_{1}\leq s\leq t$.
Proof.
Let us denote the change points of $\theta$ by $t_{1}<t_{2}<\dots<t_{M}$. Without loss of generality, $s=t_{m}$ and $t=t_{m+n}$ for $m+n\leq M$. Let us now consider a permutation $\tau:[1,\dots,M]\rightarrow[1,\dots,M]$ such that $\tau(i)=i$ for $1\leq i<m$ and $m+n<i\leq M$ and define the permuted sequence of partitions $\theta_{\tau}$ as $\theta_{\tau}(t_{m})=\theta(t_{\tau(m)})$. Despite the permutation we still get the same MCbiF for $\theta$ and $\theta_{\tau}$ for parameters $s\leq t$ because
$\bigcup_{s\leq r\leq t}\,\,\bigcup_{C\in\theta(r)}\Delta C=\bigcup_{s\leq r\leq t}\,\,\bigcup_{C\in\theta_{\tau}(r)}\Delta C.$
This implies that $\mathrm{HF}_{k}(s,t)$ is the same for $\theta$ and $\theta_{\tau}$. ∎
Appendix B Additional Examples
Our first example corresponds to the sequence of partitions analysed in Fig.1.
Example 42 (3-element example).
Let $X=\{x_{1},x_{2},x_{3}\}$ and we define $\theta(0)=\hat{0}$, $\theta(1)=\{\{x_{1},x_{2}\},\{x_{3}\}\}$, $\theta(2)=\{\{x_{1}\},\{x_{2},x_{3}\}\}$, $\theta(3)=\{\{x_{1},x_{3}\},\{x_{2}\}\}$ and $\theta(4)=\hat{1}$ so that $\theta$ is coarse-graining with $M=5$ change points. This example corresponds to Fig. 1a, and the 0- and 1-dimensional Hilbert functions are provided in Fig. 1b.
Note that $\mathrm{HF}_{0}(1,2)<|\theta(2)|$ and $\mathrm{HF}_{0}(2,3)<|\theta(3)|$ indicates the presence of two triangle 0-conflicts, which are not 1-conflicts because $\mathrm{HF}_{1}(1,2)=\mathrm{HF}_{1}(2,3)=0$. See Fig. 8a for an illustration. As shown in Proposition 31, the triangle 0-conflicts violate the strong triangle inequality of the matrix of first merge times $D_{\theta}$ (Eq 10), e.g., $D_{\theta}(x_{1},x_{3})=3>\max(D_{\theta}(x_{1},x_{2}),D_{\theta}(x_{2},x_{3}))=2$.
In addition, $\mathrm{HF}_{1}(1,3)=1$ indicates the presence of a 1-conflict that arises from the higher-order inconsistencies of cluster assignments across partitions $\theta(1)$, $\theta(2)$ and $\theta(3)$. See Fig. 8b for an illustration. In particular, the equivalence relations $x_{1}\sim_{1}x_{2}$, $x_{2}\sim_{2}x_{3}$ and $x_{3}\sim_{3}x_{1}$ induce a 1-cycle $z=[x_{1},x_{2}]+[x_{2},x_{3}]+[x_{3},x_{1}]\in Z_{1}(K^{1,3})$ and due to the lack of transitivity on the interval $[1,3]$, the 1-cycle $z$ is also non-bounding, yielding a 1-conflict. The 1-conflict then gets resolved at $t=4$ because $\theta(4)=\hat{1}$ restores transitivity on the interval $[1,4]$. See Fig. 8c for an illustration.
Figure 8: (a) Illustration of a triangle 0-conflict that violates the strong triangle inequality of the matrix of first merge times $D_{\theta,s}$ (Eq 10), (b) a 1-conflict and (c) three elements that are in no conflict due to global transitivity. If we choose $r_{1}=1,r_{2}=2,r_{3}=3$ and $r=4$, the conflicts depicted here correspond to the conflicts in Example 42.
Example 43 (4-element example).
We now consider the more complex case of a 4-element set $X=\{x_{1},x_{2},x_{3},x_{4}\}$. Let us start with $\theta(0)=\hat{0}$ and append in sequence the 6 distinct partitions that contain two singletons and one cluster of size 2, i.e., $\theta(1)=\{\{x_{1},x_{2}\},\{x_{3}\},\{x_{4}\}\}$, $\theta(2)=\{\{x_{1}\},\{x_{2},x_{3}\},\{x_{4}\}\}$, $\theta(3)=\{\{x_{1}\},\{x_{2}\},\{x_{3},x_{4}\}\}$, $\theta(4)=\{\{x_{1},x_{3}\},\{x_{2}\},\{x_{4}\}\}$, $\theta(5)=\{\{x_{1},x_{4}\},\{x_{2}\},\{x_{3}\}\}$ and $\theta(6)=\{\{x_{1}\},\{x_{2},x_{4}\},\{x_{3}\}\}$. Finally, we append consecutively three partitions, each of which contains a cluster of size 3, i.e., $\theta(7)=\{\{x_{1},x_{2},x_{3}\},\{x_{4}\}\}$, $\theta(8)=\{\{x_{1}\},\{x_{2},x_{3},x_{4}\}\}$ and $\theta(9)=\{\{x_{1},x_{3},x_{4}\},\{x_{2}\}\}$. $\theta$ is a coarse-graining, non-hierarchical sequence with $M=10$ change points. See Fig. 9 for a Sankey diagram of $\theta$.
Figure 9: Hilbert functions and Sankey diagram for the sequence of partitions $\theta$ defined in Example 43.
To analyse the topological autocorrelation of $\theta$, we compute the MCbiF Hilbert functions $\mathrm{HF}_{k}$ for dimensions $k=0,1$ (see Fig. 9). We observe that $\mathrm{HF}_{0}(s,s+3)=1<\min_{r\in[s,s+3]}|\theta(r)|$ for all $0\leq s\leq 8$, which implies that the hierarchy of $\theta$ is broken after no-less than three steps in the sequence when starting at scale $s$. Moreover, we can detect that $\theta$ is non-nested and has higher-order cluster inconsistencies because 1-conflicts emerge at scales $t=4,5,6$, as indicated by non-zero values in $\mathrm{HF}_{1}$. The 1-conflicts get resolved one-by-one through the partitions that contain clusters of size 3, and at $t=9$, when the third such partition appears in $\theta$, all 1-conflicts are resolved.
Finally, we show an example that demonstrates how conditional entropy does not detect 1-conflicts in general.
Example 44 (CE cannot detect 1-conflicts).
Let $X=\{x_{1},x_{2},x_{3},x_{4}\}$ for which we consider two different sequences of partitions $\theta(t)$ and $\eta(t)$ such that $\theta(1)=\eta(1)=\{\{x_{1},x_{2}\},\{x_{3}\},\{x_{4}\}\}$, $\theta(2)=\eta(2)=\{\{x_{1}\},\{x_{2},x_{3}\},\{x_{4}\}\}$ but $\theta(3)=\{\{x_{1},x_{3}\},\{x_{2}\},\{x_{4}\}\}\neq\theta(3)=\{\{x_{1}\},\{x_{2}\},\{x_{3},x_{4}\}\}$. See Fig. 10 for a Sankey diagram representation of the two diagrams. Note that $\theta$ and $\eta$ only differ at scale $t=3$. However, this difference is crucial because a 1-conflict emerges in $\theta$ at scale $t=0$, whereas $\eta$ has only triangle 0-conflicts and no 1-conflict. Note that $\theta$ corresponds to the toy example in Fig. 1 with one additional isolated element.
Figure 10: Sankey diagrams for sequences $\theta$ and $\eta$ defined in Example 44. Note that a 1-conflict emerges in $\theta$ at scale $t=3$, but $\eta$ has no 1-conflict.
In accordance with our theoretical results developed in Section 4.1, we can use the 1-dimensional Hilbert function $\mathrm{HF}_{1}$ to detect the 1-conflict in $\theta$ and distinguish the two sequences. In particular, $\mathrm{HF}_{1}(\theta(1),\theta(3))=1$ but $\mathrm{HF}_{1}(\eta(i),\eta(j))=0$ for all $\in[1,2,3], iji\leq j$. In contrast, the conditional entropy $\mathrm{H}$ (see Eq. 15) cannot distinguish between the two sequences as they yield the same pairwise conditional entropies. In particular, $\mathrm{H}(\theta(i)|\theta(j))=\mathrm{H}(\eta(i)|\eta(j))=\frac{1}{2}\log 2$ for $i\neq j$. This demonstrates that the conditional entropy cannot detect higher-order cluster inconsistencies in sequences of partitions.
Appendix C Details on the Homology Functor
We provide additional background on simplicial homology and its functoriality, following Hatcher [2002].
Simplicial Homology.
Let $K$ be a simplicial complex defined on the finite set $X$. For a fixed field $\boldsymbol{k}$ (the RIVET software uses the finite filed $\boldsymbol{k}=\mathbbm{Z}_{2}$ [Wright and Zheng, 2020]) and for all dimensions $k\in\{0,1,...,\dim(K)\}$ we define the $\boldsymbol{k}$-vector space $C_{k}(K)$ whose elements $z$ are given by a formal sum
$z=\sum_{\begin{subarray}{c}\sigma\in K\\ \dim(\sigma)=k\end{subarray}}a_{\sigma}\sigma$ (13)
with coefficients $a_{\sigma}\in\boldsymbol{k}$, called a $k$-chain. Note that the $k$-dimensional simplices $\sigma=[x_{0},x_{1},...,x_{k}]\in K$ form a basis of $C_{k}(K)$. For a fixed total order on $X$, the boundary operator is the linear map $\partial_{k}:C_{k}\longrightarrow C_{k-1}$ defined through an alternating sum operation on the basis vectors $\sigma=[x_{0},x_{1},...,x_{k}]$ given by
$\partial_{k}(\sigma)=\sum_{i=0}^{k}(-1)^{i}[x_{0},x_{1},...,\hat{x_{i}},...,x_{k}],$
where $\hat{x_{i}}$ means that vertex $x_{i}$ is deleted from the simplex $\sigma$. The boundary operator fulfils the property $\operatorname{im}{\partial_{k+1}}\subset\ker{\partial_{k}}$. Hence, it connects the vector spaces $C_{k}$, $k\in\{0,1,...,\dim(K)\}$, through linear maps
$\dots\xrightarrow{\partial_{k+1}}C_{k}\xrightarrow{\partial_{k}}C_{k-1}\xrightarrow{\partial_{k-1}}\dots\xrightarrow{\partial_{2}}C_{1}\xrightarrow{\partial_{1}}C_{0}\xrightarrow{\partial_{0}}0,$
leading to a sequence of vector spaces called chain complex. The elements in $Z_{k}:=\ker{\partial_{k}}$ are called $k$-cycles and the elements in $B_{k}:=\operatorname{im}{\partial_{k+1}}$ are called $k$-boundaries. Finally, the $k$-th homology group $H_{k}$ is defined as the quotient of vector spaces
$H_{k}:=Z_{k}/B_{k},$ (14)
whose elements are equivalence classes $[z]$ of $k$-cycles $z\in Z_{k}$. Each equivalence class $z]\neq 0$ corresponds to a generator of non-bounding cycles, i.e., $k$-cycles that are not the $k$-boundaries of $k+1$-dimensional simplices. This captures connected components at dimension $k=0$, holes at $k=1$ and voids at $k=2$.
Functoriality of $H_{k}$.
For fixed $k$, $H_{k}$ can be considered as a functor $H_{k}:\mathbf{Top}\rightarrow\mathbf{Vect}$, where $\mathbf{Top}$ denotes the category of topological spaces whose morphisms are continuous maps and $\mathbf{Vect}$ the category of vector spaces whose morphisms are linear maps. In particular, each topological space $K$ is sent to a vector space $H_{k}(K)$ and a continuous map $g:K\rightarrow K^{\prime}$ is sent to a linear map $H_{k}(g):H_{k}(K)\rightarrow H_{k}(K^{\prime})$ such that compositions of morphisms are preserved, i.e., $H_{k}(g\circ f)=H_{k}(g)\circ H_{k}(f)$ for two continuous maps $f$ and $g$.
Appendix D Details on Zigzag Persistence
We provide background on zigzag persistence, which was first introduced by Carlsson and de Silva [2010]. For additional details, see Dey and Wang [2022].
Zigzag Filtration.
Let $t_{1}<\dots<t_{M}$ be a sequence of real-valued parameter values. For simplicity we assume $t_{m}=m$ for $m=1,\dots,M$. Let $K^{m}$ be a simplicial complex defined on the set $X$ for every $m=1,\dots,M$. If either $K^{m}\subseteq K^{{m+1}}$ or $K^{{m+1}}\subseteq K^{{m}}$, for all $m=1,\dots,M$, we call the following diagram a zigzag filtration:
$K^{1}\leftrightarrow K^{2}\leftrightarrow\dots\leftrightarrow K^{M-1}\leftrightarrow K^{M},$
where $K^{m}\leftrightarrow K^{m+1}$ is either a forward inclusion $K^{m}\hookrightarrow K^{m+1}$ or a backward inclusion $K^{m}\hookleftarrow K^{m+1}$. While forward inclusion corresponds to simplex addition, backward inclusion can be interpreted as simplex deletion.
Zigzag Persistence.
Applying the homology functor $H_{k}$ to the zigzag filtration leads to a so called zigzag persistence module given by:
$H_{k}(K^{1})\leftrightarrow H_{k}(K^{2})\leftrightarrow\dots\leftrightarrow H_{k}(K^{M-1})\leftrightarrow H_{k}(K^{M}),$
where $H_{k}(K^{m})\leftrightarrow H_{k}(K^{m+1})$ is either a forward or backward linear map. Using quiver theory, it can be shown that a zigzag persistence module has a unique interval decomposition that provides a barcode as a simple invariant.
Appendix E Details on Information-based Baseline Methods
Information-based measures can be used to compare arbitrary pairs of partitions in the sequence $\theta$ [Meilă, 2007]. Assuming a uniform distribution on $X$, the conditional probability distribution of $\theta(t)=\{C_{1},\dots,C_{n}\}$ given $\theta(s)=\{C_{1}^{\prime},\dots,C_{m}^{\prime}\}$ is:
$P_{t|s}[i|j]=\frac{|C_{i}\cap C_{j}^{\prime}|}{|C_{j}^{\prime}|},$
and the joint probability $P_{s,t}[i,j]$ is defined similarly. The conditional entropy (CE) $\mathrm{H}(t|s)$ is then given by the expected Shannon information:
$\mathrm{H}(t|s)=-\sum_{i=1}^{|\theta(t)|}\sum_{j=1}^{|\theta(s)|}P_{s,t}[i,j]\log(P_{t|s}[i|j])$ (15)
It measures how much information about $\theta(t)$ we gain by knowing $\theta(s)$. If $\theta(s)\leq\theta(t)$ there is no information gain and $\mathrm{H}(t|s)=0$. We denote the conditional entropy matrix $\mathrm{CE}_{s,t}=H(t|s)$. Furthermore, we can compute the variation of information (VI) $\mathrm{VI}(s,t)=\mathrm{H}(s|t)+\mathrm{H}(t|s)$, which is a metric. Both $\mathrm{CE}$ and $\mathrm{VI}$ are bounded by $\log N$.
Extending information-based measures for the analysis and comparison of more than two partitions is non-trivial. However, the pairwise comparisons can be summarised with the consensus index (CI) [Vinh et al., 2010] which can be computed as the average $\mathrm{VI}$:
$\mathrm{CI}(\theta):=\frac{\sum_{i=1,i<j}^{M}\mathrm{VI}(t_{i},t_{j})}{M(M-1)/2}$ (16)
Appendix F Details on Experiments
F.1 Regression Task
Figure 11 shows the correlation between the minimal crossing number $y=\overline{\kappa}(\theta)$ (Eq. 3) and information- and MCbiF-based summary statistics. In addition to the results already described in the main text, we also observe that the correlation between $\mathrm{CI}$ and $\bar{c}_{0}$ (r=0.32r=-0.32 for $N=5$, r=0.48r=-0.48 for $N=10$) is stronger than with $\bar{c}_{1}$ (r=0.12r=-0.12 for $N=5$, r=0.34r=-0.34 for $N=10$). This can be explained by the fact that $\mathrm{CI}$ and $\bar{c}_{0}$ can both be computed from pairwise interactions of clusters in contrast to $\bar{c}_{1}$, see Section 5. Furthermore, we observe a strong correlation between $\bar{c}_{0}$ and $\bar{c}_{1}$ ($r=0.52$ for $N=5$ and $r=0.43$ for $N=10$) because of the dependencies between 0- and 1-conflicts, see Section 4.1
Figure 11: Pearson correlation ($r$) between crossing number $y$, information-based consensus index $\overline{\mathrm{CI}}$ and MCbiF-based conflict measures $\bar{c}_{0}$ and $\bar{c}_{1}$ for $N=5$ and $N=10$.
Note that we can consider our information- and MCbiF-based feature maps as $M\times M$ greyscale images, where $\mathrm{HF}_{0}$ and $\mathrm{HF}_{1}$ are symmetric and $\mathrm{CE}$ is asymmetric. The raw label encoding of $\theta$ is also similarly interpreted as an $N\times M$ greyscale image. For our regression task, we train a simple CNN [LeCun and Bengio, 1998] with one convolution and max-pool layer and one fully connected layer and also a simple MLP [Bishop, 2006] with one or two hidden layers and dropout [Srivastava et al., 2014]. For each feature map (or their combinations) separately, we perform hyperparameter optimisation for the number of filters (ranging from 2 to 6) and kernel size (chosen as 4, 8, 16, 32 or 64) in the CNN and the number of nodes (chosen as 4, 8, 16, 32, 64, 128 or 256), number of layers (1 or 2) and dropout rate (chosen as 0.00, 0.25 or 0.50) in the MLP. We use the Adam optimiser [Kingma and Ba, 2017] with learning rate chosen as 0.01, 0.005, 0.001, 0.0005 or 0.0001 for training.
We perform a full grid search of the hyperparameter space for the three different models and the different feature maps (or their combinations). We used the train split of our data for training and the validation split for evaluation and hyperparameter selection. Below, we detail the hyperparameters for the best MCbiF- and CE-based models, which were chosen according to the performance on the validation split.
  • Optimal model for $\mathrm{HF}_{0}$ & $\mathrm{HF}_{1}$ at $N=5$: CNN with 4 filters, kernel size 3, and learning rate 0.001.
  • Optimal model for $\mathrm{CE}$ at $N=5$: CNN with 8 filters, kernel size 2, and learning rate 0.005.
  • Optimal model for $\mathrm{HF}_{0}$ & $\mathrm{HF}_{1}$ at $N=10$: LR.
  • Optimal model for $\mathrm{CE}$ at $N=10$: MLP with a single layer of 256 nodes, no dropout and a learning rate of 0.001.
We present the train $R^{2}$ scores for the optimised LR, CNN and MLP models trained on the different features in Tables 3. The test $R^{2}$ scores are presented in Table 1 in the main text.
Table 3: Train $R^{2}$ scores of LR, CNN and MLP models trained on different features for $N=5$ and $N=10$.
NN Method Raw label encoding $\mathrm{HF}_{0}$ $\mathrm{HF}_{1}$ $\mathrm{HF}_{0}$ & $\mathrm{HF}_{1}$ $\mathrm{CE}$
5 LR 0.005 0.163 0.493 0.550 0.409
CNN 0.000 0.170 0.509 0.562 0.515
MLP 0.006 0.160 0.499 0.547 0.439
10 LR 0.013 0.230 0.456 0.522 0.464
CNN 0.009 0.220 0.456 0.519 0.476
MLP 0.003 0.218 0.453 0.515 0.468
F.2 Classification Task
Details on Synthetic Data.
We generate order-preserving ($y=0$) sequences $\theta\in\Pi_{N}^{M}$ through the following scheme: Let us assume that we have a total order $X=\{x_{1},\dots,x_{N}\}$ given by the element labels, i.e., $x_{i}<x_{j}$ if $i<j$. We construct each $\theta(t_{m})$, $m=0,\dots,M-1$, by cutting $X$ into clusters of the form $C=\{x_{i},x_{i+1},\dots,x_{i+n}\}$. It is easy to verify that $\theta$ is indeed order-preserving. We adapt this scheme to generate sequences $\theta\in\Pi_{N}^{M}$ that are non-order-preserving ($y=0$): Again, we start by constructing each sequence $\theta(t_{m})$ through cutting the ordered set $X$ as before. Additionally, with probability $p=0.1$, we swap the cluster assignments in $\theta(t_{m})$ for two arbitrary elements $x,y\in X$. If $N$ and $M$ are large enough, the so-generated sequence $\theta$ is almost surely non-order-preserving. We chose $N=500$ and $M=30$ to demonstrate the scalability of the MCbiF method.
The number of clusters of all our generated sequences of partitions $\theta\in\Pi_{N}^{M}$ for both classes is decreasing linearly, see Fig. 12 (a). Moreover, the average number of swaps for sequences with $y=1$ is 2.98 for our choice of $p=0.1$, see Fig. 12 (b).
Figure 12: Classification task: Histogram of the number of swaps in non-order-preserving sequences $\theta$ (class $y=1$). See text for the scheme to introduce random swaps in the node labels as a means to break order-preservation.
F.3 Application to Real-World Temporal Data
Data Preprocessing.
The temporal sequences of partitions computed by Bovet et al. [2022] are available at: https://dataverse.harvard.edu/file.xhtml?fileId=5657692. We restricted the partitions to the $N=281$ mice that were present throughout the full study period to ensure well-defined sequences of partitions, and considered the first nine temporal resolution values $\tau_{i}$, $i=1,\dots,9$, since $\theta_{\tau_{10}}$ is an outlier. Note that the sequences tend to be fine-graining, see Fig. 13.
Figure 13: Number of clusters over weeks $t$ for different temporal resolutions $\tau$, where larger values of $\tau$ produce a higher number of clusters because of the increased temporal resolution.
Figure 14: Hilbert distance between forward and backward sequences $\theta_{\tau}^{f}$ and $\theta_{\tau}^{b}$ for different temporal resolutions $\tau$.
Time Reversibility.
In the main text, we restricted our analysis to the so-called forward Flow Stability sequences of partitions. However, by reversing time direction, Bovet et al. [2022] computed a second set of backward sequences. For each temporal resolution $\tau_{i}$, we thus get a forward and backward sequence denoted by $\theta_{\tau_{i}}^{f}$ and $\theta_{\tau_{i}}^{b}$, respectively. Here we use the MCbiF to compare the forward and backward sequences of partitions for different $\tau_{i}$ and we compute the Hilbert distance $\parallel HF_{k}(\theta_{\tau_{i}}^{f})-HF_{k}(\theta_{\tau_{i}}^{b})\parallel_{2}$ for $k=0,1$, see Fig. 14.
We observe that the Hilbert distance between forward and backward sequences is high for $\tau_{2}$ because the large-scale group structure changes significantly over the study period, so that the temporal flows at low resolution $\tau_{2}$ are not reversible. In contrast, the Hilbert distance between forward and backward sequences is low for $\tau_{8}$ because the underlying social groups are more stable over the study period, leading to increased time reversibility at the high temporal resolution.